### v1

# Selecting informative conformal prediction sets with false coverage rate control

Abstract
In supervised learning, including regression and classification, conformal methods provide prediction sets for the outcome/label with finite sample coverage for any machine learning predictors. We consider here the case where such prediction sets come after a selection process. The selection process requires that the selected prediction sets be ‘informative’ in a well defined sense. We consider both the classification and regression settings where the analyst may consider as informative only the sample with prediction label sets or prediction intervals small enough, excluding null values, or obeying other appropriate ‘monotone’ constraints. While this covers many settings of possible interest in various applications, we develop a unified framework for building such informative conformal prediction sets while controlling the false coverage rate (FCR) on the selected sample. While conformal prediction sets after selection have been the focus of much recent literature in the field, the new introduced procedures, called InfoSP and InfoSCOP, are to our knowledge the first ones providing FCR control for informative prediction sets. We show the usefulness of our resulting procedures on real and simulated data.
Keywords: conformal prediction, selective inference, false coverage rate, label shift, classification, regression, prediction interval.
1 Introduction
1.1 Background
In modern data analysis, machine learning algorithms are often used to make predictions and a main challenge is to measure the uncertainty of such methods. Conformal inference offers an elegant solution to this problem, by providing prediction sets that provably cover the true value with high probability, for any sample size, any predictive algorithm, and any distribution of the data Vovk et al., (2005). We consider the following classic split/inductive conformal prediction setting Papadopoulos et al., (2002); Vovk et al., (2005); Lei et al., (2014). Let $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ be a random vector with unknown distribution $P_{XY}$. Typically, $\mathcal{X}$ is a subset of $\mathbb{R}^{d}$ (real valued covariates), and $\mathcal{Y}=[K]$ (classification of among $K\geq 2$ classes) or $\mathcal{Y}=\mathbb{R}$ (regression, with a real valued outcome) [^1] Our theory can be applied for more general observation spaces (e.g., $\mathcal{Y}=\mathbb{R}^{d^{\prime}}$) but we consider the most common settings $\mathcal{Y}\in\{[K],\mathbb{R}\}$ for simplicity.. In this setting, there are two independent samples of points $(X_{i},Y_{i})$: the calibration sample $\{(X_{j},Y_{j}),j\in[n]\}$; and the test sample $\{(X_{n+i},Y_{n+i}),i\in[m]\}$.
While all measurements are observed in the calibration sample, only the $X_{i}$’s are observed in the test sample and the aim is to provide prediction sets for the unobserved $(Y_{n+i},i\in[m])$. We denote the prediction set for $Y_{n+i}$ by $\mathcal{C}_{n+i}$. It is a subset of $[K]$ for classification and a subset (often an interval) of $\mathbb{R}$ for regression. The classic conformal prediction set for $Y_{n+i}$, denoted by $\mathcal{C}^{\alpha}_{n+i}$, is function of $\{(X_{j},Y_{j}),j\in[n]\},X_{n+i}$, and $\alpha$ (Sadinle et al.,, 2019; Lei et al.,, 2018) and has the following coverage guarantee, assuming that $(X_{n+i},Y_{n+i})$ and $(X_{j},Y_{j}),j\in[n]$, are exchangeable pairs of observations from any distribution $P_{XY}$:
$\forall\alpha\in(0,1),\>\>\mathbb{P}\left(Y_{n+i}\in\mathcal{C}^{\alpha}_{n+i}\right)\geq 1-\alpha.$ (1)
In practice, the size of the test sample $m$ is often large, encompassing hundreds or thousands of unlabeled examples. Inferring on all of them is unnecessary or inefficient in many applications (Jin and Candes,, 2023; Bao et al.,, 2024). For example, in classification, if each image belongs to one of $[K]$ categories, the analyst is not interested in the examples where $\mathcal{C}^{\alpha}_{n+i}=[K]$. Thus, it is natural to assume that a subset of individuals will be selected. However, reporting their prediction sets constructed to have at least $1-\alpha$ confidence is problematic, since conditional on being selected, the coverage may be much smaller (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013).
1.2 Informative prediction sets
Our focus is on the common setting where the analyst is only interested in reporting “interesting” or “informative” prediction sets, i.e., that cover only part of the $\mathcal{Y}$ space in a well defined sense to the analyst, which definition depends on the specific context. For concreteness, we start by providing typical examples of what can be the pre-specified collection $\mathcal{I}$ of informative subsets of $\mathcal{Y}$.
Example 1.1 (Informative prediction sets in regression, $\mathcal{Y}=\mathbb{R}$).
  1. 1.
    Intervals excluding a range of values: =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap\mathcal{Y}_{0}=\emptyset\}$ for some subset of null values $\mathcal{Y}_{0}\subset\mathcal{Y}$ that are considered as uninteresting for the user. The choice $\mathcal{Y}_{0}=(-\infty,y_{0}]$ is related to the selection proposed in Jin and Candes, (2023), for which a “normal” value for the outcome is a value below $y_{0}$.
  2. 2.
    Length-restricted intervals: =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>|I|\leq 2\lambda_{0}\}$ for some size $\lambda_{0}>0$. It is useful if the user only wants to report prediction intervals that are accurate enough.
Example 1.2 (Informative prediction sets in classification, $\mathcal{Y}=[K]$).
  1. 1.
    Excluding one class: =$\mathcal{I}=\{C\subset[K]\>:\>y_{0}\notin C\}$ for some null class $y_{0}\in[K]$. It is suitable when the user does not want to report prediction sets for individuals that are in class $y_{0}$. This can be extended to excluding several classes: =$\mathcal{I}=\{C\subset[K]\>:\>C\cap\mathcal{Y}_{0}=\emptyset\}$ for some label set $\mathcal{Y}_{0}\subset[K]$.
  2. 2.
    Non-trivial classification: =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq K-1\}$. It is appropriate when the user wants a label set that is minimally informative. More generally, at most $k_{0}$-sized classification can be considered with =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq k_{0}\}$.
A common target error guarantee is that the inference on at most $\alpha$ examples among the selected is expected to be false. This is a classical error criterion in the selective inference literature, (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013). It has been used, e.g., for novelty detection (Bates et al.,, 2021; Marandon et al.,, 2024), for classification (Rava et al.,, 2021; Zhao and Su,, 2023; Jin and Candes,, 2023), for regression (Bao et al.,, 2024), and for unsupervised clustering (Mary-Huard et al.,, 2022; Marandon et al.,, 2022). For selecting prediction sets, the target error guarantee is thus that at most $\alpha$ examples among the selected are expected to have prediction sets that do not cover their true outcome value. The false coverage proportion (FCP) for the procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is defined as the proportion of non-covered examples in the selected set $\mathcal{S}$:
$\mathrm{FCP}(\mathcal{R},Y)=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}_{n+i}\}}}{1\vee|\mathcal{S}|},$ (2)
and the target error is simply its expectation, which we refer to as the false coverage rate (FCR) as in Bao et al., (2024):
$\mathrm{FCR}(\mathcal{R})$ $=\mathbb{E}[\mathrm{FCP}(\mathcal{R},Y)].$ (3)
In this work, we consider two popular models for generating the samples [^2] In both models, the independence assumption can be relaxed. In the iid model, it is enough that all $n+m$ samples are exchangeable. In the class conditional model, it is enough that the subset of $[n+m]$ of samples from the same class is exchangeable, for all classes in $[K]$. See § A for more details.:
  • the iid model (both for regression and classification): the variables $(X_{i},Y_{i})\sim P_{XY}$, $i\in[n+m]$, are all independent and identically distributed (iid). This is the standard assumption classically used for conformal prediction (Vovk et al.,, 2005). The parameter of the model is in that case $P_{XY}$;
  • the class-conditional model (only the classification setting): all the labels $(Y_{i},i\in[n+m])$ are deterministic and the covariates $(X_{i},i\in[n+m])$ are mutually independent with $X_{j}\sim P_{X|Y=Y_{j}}$. It relaxes the exchangeable model assumptions of iid model, by only requiring that the distribution within each class is the same for the test and calibration sample (Sadinle et al.,, 2019; Ding et al.,, 2023). The target FCR is then conditional on the labels. The parameters of the model are then given by $(P_{X|Y=k})_{k\in[K]}$ and $(Y_{i},i\in[n+m])$.
1.3 Contributions
We briefly summarize here the contributions of our work.
  • We first introduce a new method, called InfoSP (Informative selective prediction set), that selects only informative prediction sets with a level $\alpha$ FCR guarantee on the selected (§ 3.1). Formally, this means that we achieve both $\mathrm{FCR}(\mathcal{R})\leq\alpha$ and $\forall i\in\mathcal{S}$, $\mathcal{C}_{n+i}\in\mathcal{I}$, for $\mathcal{I}$ being the collection of informative subsets. The FCR control of InfoSP is established both in the iid model and class-conditional model (see Theorem 3.1).
  • We introduce a second procedure, called InfoSCOP (Informative selective conditional prediction set), that has the same theoretical properties as InfoSP in the iid model (§ 3.2) and that relies on an initial selection step that is aimed at eliminating (at least some of) the examples for which informative prediction sets cannot be constructed. Further selection then takes place in order to ensure that all reported prediction sets are informative. While the pre-processing step is inspired by Bao et al., (2024), the theoretical framework in Bao et al., (2024) precludes this type of selection (see § 1.4 for more details).
  • Our main theoretical FCR control guarantees come from a single general theorem in § A.1 that generalizes the theoretical results in Benjamini and Yekutieli, (2005); Benjamini and Bogomolov, (2013), to include more general classes of $p$-values and selection rules. This goes in the direction of unifying theory for selective conformal prediction sets.
  • We optimize the analysis pipeline for common informative selection rules in § 4 (regression) and § 5-§ 7 (classification), while providing additional theoretical results and appropriate numerical experiments.
To provide an intuition for our approach, InfoSP is illustrated in Figure 1 for the regression case and in Figure 2 for the classification case (for $K=3$ classes). Left panels display a naive method reporting the marginal classical conformal prediction sets $\mathcal{C}^{\alpha}_{n+i}$ in (1) for all those that are informative, that is, such that $\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}$, without further correction. As expected, this method inflates the FCR values above $\alpha$ and thus does not provide the desired FCR control. The new procedure InfoSP is displayed in the right panels: we can see that only informative prediction intervals are selected while FCR is maintained close to $\alpha$, thanks to a suitable correction in the prediction sets.
Classical Conformal InfoSP
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Informative prediction intervals in regression (iid setting). First row: when excluding $[a,b]$ (Example 1.1 item 1) in an homoscedastic Gaussian regression model with perfect variance prediction. Second row: length-restricted (Example 1.1 item 2) in an heteroscedastic Gaussian regression model with perfect mean prediction. $\alpha=10\%$. These panels are extracted from Figures 3 and 4, see therein for more details.
Classical Conformal InfoSP
(FCR = 0.160) (FCR = 0.099)
Refer to caption Refer to caption
Figure 2: Informative prediction sets in classification for CIFAR-10 dataset, restricted to the $K=3$ classes “bird”, “cat”, and “dog” classes (iid setting). Informative prediction subsets are those of size smaller than $K-1=2$ (i.e., non-trivial, Example 1.2 item 1). Selection by InfoSP are framed in red (right panel). $\alpha=10\%$. See § 7 for more details.
1.4 Connections to existing works
We describe here inspiring works that are the most connected to ours. All these works assume the iid model.
Zhao and Su, (2023) suggested procedures for average error control in multi-class classification, so $\mathcal{C}_{n+i}$ are singletons. Since their procedure only reports a single class for each selected example, ambiguous examples will not be selected. However, in most classification tasks, there are examples whose true class is difficult to determine, yet it is possible to narrow down the possible set of classes (Sadinle et al.,, 2019). We suggest procedures that produce $\mathcal{C}_{n+i}$ that are not necessarily singletons for $K>2$. For $K=2$, their procedure coincides with an instance in our framework, see details in § 5.4. However, for $K>2$, while our suggestion as well as their suggestion provides level $\alpha$ FCR control, we select more examples, and although the prediction sets may be at a coarser resolution than singletons, they are still informative since they narrow down the possible set of classes.
Jin and Candes, (2023) addressed the problem of discovering outcomes with values above a threshold. So $\mathcal{C}_{n+i}$ is of the form $(c_{i},\infty)$ for predefined $(c_{i})_{i\in[m]}$. They cast the problem as that of testing the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$. In § E.2, it is demonstrated that by defining $Y_{n+i}\leq c_{i}$ as uninformative, we can complement the discoveries of Jin and Candes, (2023) with one-sided prediction intervals, while providing the same false discovery rate guarantee on the selected. Moreover, we show how to obtain two-sided prediction intervals for the informative examples in § 4.1.
Bao et al., (2024) considered the regression framework. Their first result (Proposition 1 therein) is to prove that for selection rules that do not depend on the calibration sample, classic conformal prediction intervals at level $\alpha|\mathcal{S}|/m$ for the $\mathcal{S}$ selected examples (i.e., the correction factor $|\mathcal{S}|/m$ for selection suggested in Benjamini and Yekutieli,, 2005), provide level $\alpha$ FCR control. For our purpose of informative selection, this result is not useful because informative selection involves all conformal $p$-values and therefore involves the calibration sample in a specific way. We need a different set of assumptions that are detailed in our novel Theorem A.1.
Bao et al., (2024) further argue that the resulting prediction intervals, $\big{(}\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}\big{)}_{i\in\mathcal{S}}$ are too wide. They suggest a novel approach that performs selection on both the calibration set and test set, and then constructs $\alpha$ level conformal prediction intervals for the selected test candidates using the conditional empirical distribution obtained by the post-calibration set. For exchangeable selection rules, they show that the FCR is controlled at level $\alpha$. Their selection process cannot guarantee that all the constructed prediction intervals are of interest to the analyst. For example, for predicting the affinity of drug-target pairs, the analyst may not be interested in pairs with affinity below, say, $y_{0}$ (the case in item 1 of Example 1.1). Using the novel approach of Bao et al., (2024), prediction intervals will be constructed following the selection of calibration and test examples for which the predicted affinity from the machine learning algorithm is above a selection threshold. They require that the selection procedure be a thresholding procedure of the scores $S_{Y_{j}}(X_{j})$ with a threshold $\tau$ that is either independent of the calibration sample (their Proposition 1), or exchangeable with respect to both calibration and test samples (their Theorem 1). Some of these prediction intervals may include $y_{0}$, and thus be useless for the analyst (an illustration is given in § E.3 in SM, see Figure 12 therein). However, it is not possible to additionally select only the examples with prediction intervals above $y_{0}$, since after performing this additional selection, the prediction intervals of Bao et al., (2024) on the selected no longer have an $\alpha$ level FCR guarantee. Our procedures for regression thus complement the work of Bao et al., (2024) when the focus is that all the prediction intervals eventually constructed are informative.
In a very recent work [^3] This work appeared when we were in the final stage of writing, our work has been done independently., Jin and Ren, (2024) generalize the work of Bao et al., (2024), by considering more general selection rules for finite-sample exact coverage conditional on the unit being selected. Their conditional guarantee is achieved by a careful swapping argument which identifies for each selection rule, the appropriate subset of the calibration examples for each example from the test sample. Their conditional error guarantee implies FCR control under some conditions. They provide various examples where selection takes place and prediction sets on the selected are of interest. Even though their selection may be based on some notion of informativeness, like in Bao et al., (2024) the final prediction sets that are constructed can violate this notion. For example, after selecting by a multiple testing procedure on the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$ in the setting of Jin and Candes, (2023), their prediction sets may include the $c_{i}$’s for some of the discoveries. We suggest procedures where selection and construction of prediction sets are inseparable, since we require that each selected prediction set be informative (along with the requirement that FCR $\leq\alpha$).
Notation
Expectations and probabilities are denoted for the iid model with $\mathbb{E}_{(X,Y)\sim P_{XY}}(\cdot)$ and $\mathbb{P}_{(X,Y)\sim P_{XY}}(\cdot)$, and for the class-conditional model with $\mathbb{E}_{X\sim P_{X|Y}}(\cdot)$ and $\mathbb{P}_{X\sim P_{X|Y}}(\cdot)$, respectively. If the data generation process is clear from the context, or if the expression is relevant for both models, then the subscript is omitted. In addition, $A\subset B$ means that the set $A$ is included in the set $B$ with a possible equality. Finally, for any subset $C\subset\mathbb{R}$, $|C|$ denotes the cardinality of $C$ if the set $C$ is finite, and the Lebesgue measure of $C$ if $C$ is an interval.
2 Preliminaries
2.1 Conformal prediction sets
The classical conformal prediction set for $Y_{n+i}$ is given by
$\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})=\{y\in\mathcal{Y}\>:\>{p}^{(y)}_{i}>\alpha\},\>\>\>i\in[m],$ (4)
where $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m]\}$ is a collection of conformal $p$-values satisfying the following super-uniform guarantee:
$\mathbb{P}(p^{(Y_{n+i})}_{i}\leq\alpha)\leq\alpha,\>\>\>i\in[m],$ (5)
where the above probability is compute either in the iid ($Y_{n+i}$ random) or conditional ($Y_{n+i}$ fixed) model. Super-uniformity (5) implies directly that $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) provides $1-\alpha$ coverage for $Y_{n+i}$, that is,
$\mathbb{P}(Y_{n+i}\in\mathcal{C}^{\alpha}_{n+i}(\mathbf{p}))\geq 1-\alpha,\>\>\>i\in[m],\>\>\>\alpha\in(0,1),$ (6)
which is generally referred to as marginal coverage.
The $p$-value family is built from the calibration and test sample by using non-conformity score functions $S_{y}:x\in\mathbb{R}^{d}\mapsto\mathbb{R}$, $y\in\mathcal{Y}$, measuring the inadequacy between $y$ and the prediction at point $x$. Importantly, we follow a split/inductive conformal approach, where the score functions have been computed from an independent training data sample so that they can be considered as fixed here [^4] Note that we can relax slightly this assumption: our theory also allows this function to depend on the calibration plus test samples in an exchangeable way, see Assumption 6 in the SM. (and all probabilities/expectations are taken conditional on that training sample).
Assumption 1.
The score functions $S_{y}(\cdot)$, $y\in\mathcal{Y}$, have been computed from an independent training data sample and the computed scores $(S_{Y_{i}}(X_{i}),i\in[n+m])$ have no ties almost surely. Moreover, when $\mathcal{Y}=\mathbb{R}$, the score function is regular in the following sense: for every $x\in\mathbb{R}^{d}$, the function $y\in\mathbb{R}\mapsto S_{y}(x)\in\mathbb{R}$ is right continuous with left limits.
Assumption 1 is a very mild assumption, which is typically satisfied. For instance, for the regression case, a classical choice is the locally weighted residual function $S_{y}(x)=|y-\mu(x)|/\sigma(x)$ where $\mu(x)\in\mathcal{Y}$ is the predicted outcome at point $x$ and $\sigma(x)$ is the predicted standard deviation of $Y$ given $X=x$ (Lei et al.,, 2018). Another common example is the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$ where $q_{\beta}(x)$ is the predicted $\beta$-quantile of $Y$ given $X=x$, which corresponds to the so-called conformalized quantile regression (Romano et al.,, 2019; Sesia and Romano,, 2021) (for some prespecified $10<\beta_{0}<\beta_{1}<$ 1). More generally, we refer the reader to Gupta et al., (2022) for a general framework giving rise to a large class of score functions. In the classification case, the typical score is the residual function $S_{y}(x)=1-\pi_{y}(x)$ where $\pi_{y}(x)$ is an estimator of the probability to generate label $y$ at point $x$.
Formally, the $p$-value family is given as follows:
  • full-calibrated $p$-values: $\mathbf{\overline{p}}=(\overline{p}^{(y)}_{i},i\in[m],y\in\mathcal{Y})$, both for the regression and classification cases, with
    $\overline{p}^{(y)}_{i}=\frac{1}{n+1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{y}(X_{n+i})\}}\Big{)},\>\>i\in[m],y\in\mathcal{Y}.$ (7)
  • class-calibrated $p$-values: $\mathbf{{\widetilde{p}}}=(\tilde{p}^{(y)}_{i},i\in[m],y\in\mathcal{Y})$, only for the classification case $\mathcal{Y}=[K]$, with
    $\tilde{p}^{(y)}_{i}=\frac{1}{|\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}|+1}\Big{(}1+\sum_{j\in\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}}{\mathbf{1}\{S_{y}(X_{j})\geq S_{y}(X_{n+i})\}}\Big{)},\>\>i\in[m],y\in\mathcal{Y},$ (8)
    where $\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}=\{j\in[n]\>:\>Y_{j}=y\}$ corresponds to the elements of the calibration sample that have label $y\in\mathcal{Y}$.
Both $p$-values $\overline{p}^{(y)}_{i}$ and $\tilde{p}^{(y)}_{i}$ are computed by examining how extreme the score $S_{y}(X_{n+i})$ is among the scores of the true labels in the calibration sample. Full-calibrated $p$-values and class-calibrated $p$-values satisfy the super-uniformity (5) in the iid model and class-conditional model, respectively, by using an exchangeability argument, see Vovk et al., (2005); Romano and Wolf, (2005); Bates et al., (2021). This means that prediction set $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) satisfies the marginal coverage (6) in each context. As a result, the false coverage rate (3) for the full selection $\mathcal{S}=[m]$ is controlled at level $\alpha$, that is,
$\mathbb{E}\left[\frac{\sum_{i\in[m]}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\}}}{m}\right]=m^{-1}\sum_{i\in[m]}\mathbb{P}(Y_{n+i}\notin\mathcal{C}^{\alpha}_{n+i}(\mathbf{p}))\leq\alpha.$ (9)
Remark 2.1.
An equivalent description of the prediction set $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) can be given as a score level set, with a threshold depending on the calibration sores. Formally, we have
$\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})=\{y\in\mathcal{Y}\>:\>p^{(y)}_{i}>\alpha\}=\{y\in\mathcal{Y}\>:\>S_{y}(X_{n+i})\leq\hat{s}_{\alpha}\}$ (10)
[…]
2.2 $\mathcal{I}$-adjusted $p$-values
Our theory relies on the following assumption.
Assumption 2.
The subset collection $\mathcal{I}$ is monotone in the following sense: for the considered $p$-value collection $\mathbf{p}$ (either $\mathbf{\overline{p}}$ or $\mathbf{{\widetilde{p}}}$), we have
  • (i)
    If a prediction set is informative, then all the prediction sets it contains are also informative, that is,
    • Classification: for all $\mathcal{C},\mathcal{C}^{\prime}\subset[K]$, if $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{C}\in\mathcal{I}$, then $\mathcal{C}^{\prime}\in\mathcal{I}$.
    • Regression: for all $\mathcal{C},\mathcal{C}^{\prime}$ intervals of $\mathbb{R}$, if $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{C}\in\mathcal{I}$, then $\mathcal{C}^{\prime}\in\mathcal{I}$.
  • (ii)
    Almost surely, the function $\alpha\in(0,1]\mapsto{\mathbf{1}\{\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}\}}\in\{0,1\}$ is right-continuous.
  • (iii)
    (Only in the regression case) For all $\alpha\in(0,1)$, almost surely, $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ is an interval of $\mathbb{R}$.
Note that Assumption 2 implies that $\alpha\in(0,1]\mapsto{\mathbf{1}\{\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}\}}\in\{0,1\}$ is both right-continuous and nondecreasing: if $\alpha\leq\alpha^{\prime}$, it follows that $\mathcal{C}^{\alpha^{\prime}}_{n+i}(\mathbf{p})\subset\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ from (4) and thus $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}$ implies that $\mathcal{C}^{\alpha^{\prime}}_{n+i}(\mathbf{p})\in\mathcal{I}$ by Assumption 2 (i) (iii).
As a result, we can define the $\mathcal{I}$-adjusted $p$-value vector by $\mathbf{q}=(q_{i})_{i\in[m]}$ where
$q_{i}=\min\{\alpha\in(0,1]\>:\>\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}\},$ (11)
with by convention $q_{i}=1$ if the set is empty. Assumption 2 can be easily checked for Examples 1.1 and 1.2, with an explicit expression for $q_{i}$’s.
Example 2.1 (Example 1.1 continued).
For regression, that is $\mathcal{Y}=\mathbb{R}$ (see § 4 for more details):
  1. 1.
    Intervals excluding $\mathcal{Y}_{0}=[a,b]$: $q_{i}=\max_{y\in[a,b]}p^{(y)}_{i}$ (if this maximum exists).
  2. 2.
    Length-restricted intervals $I$ (with $|I|\leq 2\lambda_{0}$):
    • for the locally weighted score function $S_{y}(x)=|y-\mu(x)|/\sigma(x)$,
      $q_{i}=(n+1)^{-1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})>\lambda_{0}/\sigma(X_{n+i})\}}\Big{)};$ (12)
    • for the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$,
      $q_{i}=(n+1)^{-1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})>\lambda_{0}-(q_{\beta_{1}}(X_{n+i})-q_{\beta_{0}}(X_{n+i}))/2\}}\Big{)}.$ (13)
Example 2.2 (Example 1.2 continued).
For classification, that is, $\mathcal{Y}=[K]$:
  1. 1.
    Excluding one-class: $q_{i}=p^{(y_{0})}_{i}$; or excluding several classes: $q_{i}=\max_{y\in\mathcal{Y}_{0}}p^{(y)}_{i}$.
  2. 2.
    Non trivial classification: $q_{i}=\min_{y\in[K]}p^{(y)}_{i}$; or at most $k_{0}$-sized classification: $q_{i}$ is the $(K-k_{0})$-th smallest element in the set $\{p^{(y)}_{i},y\in[K]\}$.
Example 2.3 (Combining informative subset collections).
Let two subset collections $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ that satisfy Assumption 2 with adjusted $p$-values given by $(q_{1,i})_{i\in[m]}$ and $(q_{2,i})_{i\in[m]}$, respectively. Then we can easily check that the intersected collection
:=$\mathcal{I}:=\{I_{1}\cap I_{2},I_{1}\in\mathcal{I}_{1},I_{2}\in\mathcal{I}_{2}\}$
also satisfies Assumption 2 with adjusted $p$-values given by $q_{i}=\max\{q_{1,i},q_{2,i}\}$, $i\in[m]$. This is useful to combine the constraints imposed by the informativeness. For instance, in the regression case (Examples 1.2 and 2.2), we can declare a subset as informative if it excludes a null class while it is of cardinality at most $1$ (see § 6 for a concrete application).
From its definition in (11), it follows that $q_{i}$ can be seen as a $p$-value to test whether $Y_{n+i}$ lies in an informative set or not, that is, to test
$H_{0,i}$: “$Y_{n+i}\notin\cup_{C\in\mathcal{I}}C$” versus $H_{1,i}:$$Y_{n+i}\in\cup_{C\in\mathcal{I}}C$”. (14)
In cases where being informative is linked to a particular location in $\mathcal{Y}$, this testing problem takes an especially meaningful form:
  • Excluding one class in classification: $q_{i}=p^{(y_{0})}_{i}$ tests $H_{0,i}$: “$Y_{n+i}=y_{0}$” versus $H_{1,i}:$$Y_{n+i}\neq y_{0}$”;
  • Excluding $\mathcal{Y}_{0}=[a,b]$ in regression: $q_{i}$ tests $H_{0,i}$: “$Y_{n+i}\in[a,b]$” versus $H_{1,i}:$$Y_{n+i}\notin[a,b]$”.
Note that in case where informative sets are those with small size (e.g., non-trivial classification or length-restricted intervals), we have $\cup_{C\in\mathcal{I}}C=\mathcal{Y}$, so the null hypothesis is always false, and the testing problem is not interesting.
Remark 2.2.
Assumption 2 implies that the $\mathcal{I}$-adjusted $p$-value vector $\mathbf{q}$ is a nondecreasing function of the $p$-value collection, see Lemma A.5; Assumption 2 (iii) is always satisfied up to taking as prediction sets $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ the convex hull of the set in the right-hand-side of (4). It is also satisfied without modifying $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ for any score function such that all the sets $\{y\in\mathbb{R}\>:\>S_{y}(x)\leq s\}$ are intervals, which is often the case, see Remark 2.1 and § 4.
2.3 Aim: informative selection with FCR guarantees
The general inferential task is to produce prediction sets for examples of interest in the test sample, that is, after selection. The selection process is driven by the requirement that the prediction sets be informative, and the requirement of a relevant error control.
Definition 1.
A selective prediction set procedure is a function of the observations $\{(X_{j},Y_{j}),j\in[n]\}$, $\{X_{n+i},i\in[m]\}$ of the form $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ where $\mathcal{S}$ is the selected subset of $[m]$ for which prediction sets are constructed, and $\mathcal{C}_{n+i}\subset\mathcal{Y}$ is the prediction set for $Y_{n+i},i\in\mathcal{S}$.
Definition 2.
A selective prediction set procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is said to be $\mathcal{I}$-informative (or informative) if the selection $\mathcal{S}$ is a subset of $[m]$ that imposes that $\mathcal{C}_{n+i}$ is informative, that is, $\forall i\in\mathcal{S}$, $\mathcal{C}_{n+i}\in\mathcal{I}$.
The FCR in the iid model and class-conditional model, respectively, are
$\mathrm{FCR}(\mathcal{R},P_{XY})$ $=\mathbb{E}_{(X,Y)\sim P_{XY}}[\mathrm{FCP}(\mathcal{R},Y)];$ (15)
$\mathrm{FCR}(\mathcal{R},P_{X\mid Y},Y)$ $=\mathbb{E}_{X\sim P_{X\mid Y}}[\mathrm{FCP}(\mathcal{R},Y)],$ (16)
for the FCP in (2) The FCR expression in (16) for quantifying the errors among the selected is classical in the selective inference literature (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013), since $(Y_{n+i})_{i\in[m]}$ are fixed so this is the false coverage rate on the parameters. On the other hand, in (15), we consider the false coverage rate of prediction intervals. So the false coverage rate is on random outcomes (i.e., $(Y_{n+i})_{i\in[m]}$ in our setting) rather than on parameters, which is related to the Bayes FDR criterion, see, e.g., Efron et al., (2001).
We will focus on finding selective prediction set procedures $\mathcal{R}=\mathcal{R}_{\alpha}$ with either of the two following controls:
$\sup_{P_{X,Y}}\{\mathrm{FCR}(\mathcal{R},P_{X,Y})\}\leq\alpha\>\>;$ (17)
$\sup_{P_{X\mid Y},Y}\{\mathrm{FCR}(\mathcal{R},P_{X\mid Y},Y)\}\leq\alpha\>\>.$ (18)
Obviously, the control (18) (class-conditional control, only considered for classification) is stronger than (17) (unconditional control, considered both for classification and regression).
To balance with FCR control, we also consider a notion of power that is the “resolution adjusted” correct decision number:
$\mathrm{Pow}(\mathcal{R})$ $=\mathbb{E}\left[\sum_{i\in\mathcal{S}}\frac{{\mathbf{1}\{Y_{n+i}\in\mathcal{C}_{n+i}\}}}{|\mathcal{C}_{n+i}|}\right].$ (19)
Hence, for the same selection set, a decision with a smaller covering decision set $\mathcal{C}_{n+i}$ yields higher power. Overall, we aim to maximize the power while controlling the FCR at a pre-specified level $\alpha$. This corresponds to the situation where the analyst would like to select as many examples as possible that are informative, with as narrow as possible a prediction set for the selected examples, while controlling the FCR.
Finally, considering the multiple testing problem (14), that test if $Y_{n+i}$ lies in an informative set or not, we can also quantify the error amount of a given selection procedure $\mathcal{S}$ (by itself, without quantifying the non-coverage errors of the attached prediction sets), by letting (Benjamini and Hochberg,, 1995)
$\mathrm{FDP}(\mathcal{S},Y)$ $=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\cup_{C\in\mathcal{I}}C\}}}{1\vee|\mathcal{S}|}.$ (20)
For instance, when one wants to exclude a given label set $\mathcal{Y}_{0}$ in classification/regression, we have $\mathrm{FDP}(\mathcal{S},Y)=(\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\in\mathcal{Y}_{0}\}})/(1\vee|\mathcal{S}|)$. The false discovery rates in the iid model and class-conditional model are the corresponding expectations
$\mathrm{FDR}(\mathcal{S},P_{XY})$ $=\mathbb{E}_{(X,Y)\sim P_{XY}}[\mathrm{FDP}(\mathcal{S},Y)];$ (21)
$\mathrm{FDR}(\mathcal{S},P_{X\mid Y},Y)$ $=\mathbb{E}_{X\sim P_{X\mid Y}}[\mathrm{FDP}(\mathcal{S},Y)],$ (22)
respectively. The following lemma holds.
Lemma 2.1.
For any selective prediction set procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ that is $\mathcal{I}$-informative, we have $\mathrm{FDP}(\mathcal{S},Y)\leq\mathrm{FCP}(\mathcal{R},Y).$
It comes directly from the fact that, for $i\in\mathcal{S}$, $Y_{n+i}\in\mathcal{C}_{n+i}$ implies $Y_{n+i}\in\cup_{C\in\mathcal{I}}C$ because $\mathcal{C}_{n+i}\in\mathcal{I}$. As a result, producing an informative selective prediction set procedure that controls the FCR at level $\alpha$ immediately ensures that the attached selection procedure controls the FDR at level $\alpha$ and thus comes with a relevant interpretation.
3 Main results
3.1 Informative selective prediction sets (InfoSP)
In order to have a level $\alpha$ FCR guarantee on the selected examples, it is necessary to correct the threshold $\alpha$ in $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ (4). A standard approach in the selective inference literature (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013) is to use the reduced level $\alpha|\mathcal{S}|/m$ for a selection set $\mathcal{S}$. In order for the selective prediction set procedure to be $\mathcal{I}$ informative, the selection rule needs to be carefully selected. We shall use the following basic observation: selection by a thresholding rule on the family $\mathbf{q}=(q_{i},i\in[m])$ given by (11) will result in selected examples that are $\mathcal{I}$ informative, since $\mathcal{C}^{t}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq t$ for all $t\in(0,1]$ by the definition of $q_{i}$. Combining the standard approach for FCR control with this basic observations, leads us to use the following selection thresholding rule which is necessarily $\mathcal{I}$-informative: all examples with $q_{i}$ at most
$\max\left\{t:t\leq\alpha\frac{\left(\sum_{j=1}^{m}{\mathbf{1}\{q_{j}\leq t\}}\right)}{m}\right\}.$
This is exactly the BH selection rule (Benjamini and Hochberg,, 1995) on the adjusted $p$-value vector $\mathbf{q}=(q_{i},i\in[m])$. In practice, let us recall that the BH procedure $\mathrm{BH}(\mathbf{q})$ can be obtained as
$\mathrm{BH}(\mathbf{q})=\{i\in[m]\>:\>q_{i}\leq\alpha\hat{\ell}/m\},$ (23)
where ^$\hat{\ell}=|\mathrm{BH}(\mathbf{q})|=\max\{\ell\in[m]\>:\>q_{(\ell)}\leq\alpha\ell/m\}$ (with ^$\hat{\ell}=0$ if the set is empty) and where $q_{(1)}\leq\dots\leq q_{(m)}$ denote the ordered $q_{i}$’s.
Definition 3.
The informative selective prediction set procedure (InfoSP) based on a $p$-value family $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m]\}$, is defined as $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})=(\mathcal{C}^{\alpha|\mathrm{BH}(\mathbf{q})|/m}_{n+i}(\mathbf{p}))_{i\in\mathrm{BH}(\mathbf{q})}$, that is, is given as follows:
  1. 1.
    Apply the BH procedure on the corresponding adjusted vector $\mathbf{q}=(q_{i},i\in[m])$ given by (11),(23) and select $\mathcal{S}(\mathbf{p})=\mathrm{BH}(\mathbf{q})\subset[m]$;
  2. 2.
    For each $i\in\mathcal{S}(\mathbf{p})$, consider the prediction set $\mathcal{C}^{\alpha|\mathcal{S}(\mathbf{p})|/m}_{n+i}(\mathbf{p})$ for $Y_{n+i}$, computed according to (4).
The key theoretical difficulty in proving that the FCR of InfoSP is at most $\alpha$, is that the conformal $p$-values are dependent, and that the selection step is also $p$-value-based (by contrast with Bao et al., (2024)). Thus, the error rate in (17) for the iid model and (18) for the class-conditional model may not be controlled. Bates et al., (2021); Marandon et al., (2024) showed that, for outlier detection, the special positive dependency between the conformal $p$-values is such that the BH procedure remains valid for FDR control. Their results are on a different set of conformal $p$-values, but we develop a similar result for our problem, which enables us to establish the desired error guarantee for various selection rules.
Theorem 3.1.
Consider score functions satisfying Assumption 1, an informative subset collection $\mathcal{I}$ satisfying Assumption 2 and a $p$-value collection $\mathbf{p}$ being either $\mathbf{\overline{p}}$ or $\mathbf{{\widetilde{p}}}$, then the information selective prediction set procedure $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ defined in Definition 3 controls the FCR at level $\alpha$, respectively in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) with the control (17) or the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$) with the control (18).
A proof is provided in § C.1, which relies on a more general result, see Theorem A.1 in § A. Theorem A.1 extends the FCR guarantee to more general $p$-value collections (Assumption 4) and more general selection rules (Assumption 5). In particular, it is useful to note that our result applies beyond informative selection (e.g., selection by $p$-value thresholding) and for the two $p$-value collections $\mathbf{\overline{p}}$ and $\mathbf{{\widetilde{p}}}$ under less restrictive conditions that those considered in Theorem 3.1: the independence assumption can be relaxed to an exchangeable assumption (Propositions A.2 and A.3), while the score function can take a general form that can provide an extra improvement (Assumption 6).
Applying Lemma 2.1, we obtain the following theoretical guarantee on the selection rule of InfoSP.
Corollary 3.2.
In the setting of Theorem 3.1, the attached selection rule $\mathrm{BH}(\mathbf{q})$ of $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ controls the FDR given either by (21) or (22) at level $\alpha$, respectively in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) or the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$).
Remark 3.1.
For InfoSP, we can avoid the computation of $\mathbf{q}$ by using that the BH procedure is the iterative limit of a recursion (Gao et al.,, 2023). Indeed, since $\mathcal{C}^{\alpha k/m}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq\alpha k/m$, the selection $S=\mathrm{BH}(\mathbf{q})$ can be obtained as follows:
  • Step 1: $\mathcal{S}_{1}=\{i\in[m]\>:\>\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}\}$;
  • Step $t\geq 1$: $\mathcal{S}_{t}=\{i\in\mathcal{S}_{t-1}\>:\>\mathcal{C}^{\alpha|\mathcal{S}_{t-1}|/m}_{n+i}\in\mathcal{I}\}$;
  • Consider $t_{0}$ the first $t$ where $\mathcal{S}_{t}=\mathcal{S}_{t-1}$ and let $\mathcal{S}=\mathcal{S}_{t_{0}}$.
3.2 Informative selective conditional prediction sets (InfoSCOP)
Throughout this section, we consider the iid model. It turns out that InfoSP can be too conservative in some contexts; this is manifest in the true FCR being much smaller than the nominal $\alpha$ level, see illustrations in § 4 and § 5. To avoid this, we can adapt the conditional approach of Bao et al., (2024) to our framework and start by selecting test samples and calibration samples. We would like, following the initial selection rule to have as few as possible test samples for which $\mathcal{I}$-informative prediction sets cannot be constructed. Specifically, we follow the following steps
  1. 1.
    Split the calibration sample $((X_{j},Y_{j}),j\in[n])$ into two samples $((X_{j},Y_{j}),j\in[r])$ and $((X_{j},Y_{j}),j\in[r+1,n])$ for some $r\in[n-1]$.
  2. 2.
    Apply an initial conformal selection rule $\mathcal{S}^{(0)}=\mathcal{S}^{(0)}((X_{j},Y_{j})_{j\in[r]},(X_{j})_{j\in[r+1,n+m]})\subset[r+1,n+m]$ that considers as calibration sample $((X_{j},Y_{j}),j\in[r])$ and test sample $((X_{j},Y_{j}),j\in[r+1,n+m])$.
  3. 3.
    For $i+n\in\mathcal{S}^{(0)}\cap[n+1,n+m]$, compute the conformal $p$-values using calibration set $\{(X_{j},Y_{j}),j\in\mathcal{S}^{(0)}\cap[r+1,n]\}$:
    $\overline{p}^{(0,y)}_{i}=\frac{1}{|\mathcal{S}^{(0)}\cap[r+1,n]|+1}\left(1+\sum_{j\in\mathcal{S}^{(0)}\cap[r+1,n]}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{y}(X_{n+i})\}}\right).$ (24)
Also, we assume that the initial selection $\mathcal{S}^{(0)}$ satisfies the following permutation preserving assumption.
Assumption 3.
For any permutation $\sigma$ of $[r+1,n+m]$,
$\mathcal{S}^{(0)}((X_{j},Y_{j})_{j\in[r]}),(X_{\sigma(j)})_{j\in[r+1,n+m]})=\sigma\big{(}\mathcal{S}^{(0)}((X_{j},Y_{j})_{j\in[r]},(X_{j})_{j\in[r+1,n+m]})\big{)}.$
The initial selection $\mathcal{S}^{(0)}$ is typically the result of a multiple testing procedure that is applied to the examples in $[r+1,n+m]$, that uses $p$-values computed using $((X_{j},Y_{j}),j\in[r])$ as calibration sample and $(X_{i},i\in[r+1,n+m])$ as covariate test sample, which immediately satisfies Assumption 3. We can use $\mathrm{BH}(\mathbf{q})$ as an initial selection stage (where the $q_{i}$’s are computed with the aforementioned sample split) so that selected examples are likely to correspond to $Y_{n+i}$ where an informative prediction set can be built. For excluding a null range in regression or a null class in classification, another choice is to use an appropriate BH procedure for testing that the examples from $[r+1,n+m]$ are from that null.
Definition 4.
The informative selective conditional prediction set procedure pre-processed with the initial selection rule $\mathcal{S}^{(0)}$, called InfoSCOP, is defined as the InfoSP procedure of Definition 3 applied with the pre-processed $p$-value family $\mathbf{\overline{p}}^{0}$ (24), that is,
$\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})=\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}^{0})=(\mathcal{C}^{\alpha^{0}}_{n+i}(\mathbf{\overline{p}}^{0}))_{i\in\mathrm{BH}(\mathbf{q}^{0})},$
where $\alpha^{0}=\alpha|\mathrm{BH}(\mathbf{q}^{0})|/|\mathcal{S}^{(0)}\cap[n+1,n+m]|$ and the $\mathcal{I}$-adjusted $p$-values $\mathbf{q}^{0}$ are computed via (11) from the pre-processed $p$-values $\mathbf{\overline{p}}^{0}$.
Theorem 3.3.
Consider the iid model (both for regression and classification), score functions satisfying Assumption 1, an informative subset collection $\mathcal{I}$ satisfying Assumption 2 and any initial selection rule $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies Assumption 3. Then the InfoSCOP procedure of Definition 4 is such that $\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}}))\leq\alpha$. In addition, the associated selection rule $\mathrm{BH}(\mathbf{q}^{0})$ controls the FDR (21) at level $\alpha$.
The proof is provided in § C.2.
As we will see in the next sections, while it maintains the FCR guarantee, InfoSCOP can greatly improve over InfoSP. The main reason is that adjusting for selection is cheaper after the initial selection step: by reducing the fraction of examples in the test sample for which it is not possible to construct $\mathcal{I}$ informative prediction sets, the correction term $\alpha^{0}$ is expected to be close to $\alpha$ (or not much smaller than $\alpha$). Another reason is that the selection-conditional $p$-values $\mathbf{\overline{p}}^{0}$ will be better in settings where the initial selection step tends to remove the examples from the calibration set that have large non-conformity scores, see § 4, § E and § F for such cases.
In general, the way InfoSCOP can improve over InfoSP depends on the context. For instance, for excluding a null range in classification, a null class in regression or for length restriction in regression, we show in § 4 and § 5, respectively, the great potential advantage of using the initial selection. On the other hand, we also show that for non-trivial classification, there may be no advantage of initial selection (in fact, there can be a light disadvantage since the calibration sample after initial selection is smaller, as demonstrated in § 5.4).
4 Application to regression
This section is devoted to the regression case (that is, $\mathcal{Y}=\mathbb{R}$), as already introduced in Example 1.1. Throughout the section, we work in the iid model. Illustrations for other informative selections can be found in § E.
4.1 Excluding an interval $[a,b]$ of the prediction interval
We focus here on the case where the user wants to build prediction intervals only for labels such that $Y_{n+i}<a$ or $Y_{n+i}>b$, which corresponds to excluding the interval $\mathcal{Y}_{0}=[a,b]$ of the prediction interval, where $a<b$ are two benchmark values. This corresponds to common practice where users are interested only in reporting prediction intervals for individuals with “abnormal” outcomes. Setting $a=-\infty$ recover the case where we only want to report prediction interval for examples such that $Y_{n+i}>b$, which is the selection considered in Jin and Candes, (2023). We focus on two-sided prediction intervals here (the case of one-sided prediction intervals is postponed to § E). The choice of score function defines InfoSP and entails all the desired inferential guarantees. We formalize this for the locally weighted score function in Corollary 4.1. Using other score functions is also possible, e.g., the score function that corresponds to conformal quantile regression, see Remark 4.1.
Corollary 4.1.
Consider the iid model in the regression case, consider the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ and suppose that Assumption 1 holds. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap[a,b]=\emptyset\}$ and full calibrated $p$-value collection $\mathbf{\overline{p}}$ (7):
  • (i)
    InfoSP selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ with
    $q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}},$ (25)
    where $\overline{p}^{(a)}_{i}$ and $\overline{p}^{(b)}_{i}$ are the full-calibrated $p$-values (7).
  • (ii)
    The selection $\mathcal{S}$ of InfoSP controls the FDR at level $\alpha$ in the following sense:
    $\sup_{P_{XY}}\mathbb{E}_{(X,Y)\sim P_{XY}}\left[\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\in[a,b]\}}}{1\vee|\mathcal{S}|}\right]\leq\alpha.$
  • (iii)
    The selected prediction intervals do not contain $[a,b]$ and are of the form $\mathcal{C}_{n+i}=[\mu(x)-S_{(n_{\alpha}(\mathbf{p}))}\sigma(x),\mu(x)+S_{(n_{\alpha}(\mathbf{p}))}\sigma(x)]$, where $S_{(1)}\leq\dots\leq S_{(n)}$ are the ordered calibration scores $S_{Y_{j}}(X_{j})$, $1\leq j\leq n$ (with $S_{(n+1)}=+\infty$), and $n_{\alpha}(\mathbf{p})=\lceil(1-\alpha|\mathcal{S}(\mathbf{p})|/m)(n+1)\rceil$.
  • (iv)
    These prediction intervals control the FCR at level $\alpha$ in the sense of (17).
Proof.
Point (i) comes from Example 2.1 because $\min_{y\in[a,b]}S_{y}(x)$ is 00 if $\mu(x)\in[a,b]$, $S_{a}(x)$ if $\mu(x)<a$, and $S_{b}(x)$ if $\mu(x)>b$. Point (ii) comes from Lemma 2.1, (iii) from Remark 2.1 and the fact that $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$ iff $\mathcal{C}_{n+i}$ does not contain $[a,b]$. Point (iv) comes from Theorem 3.1. ∎
Hence, our method, in addition to providing an FCR control on the selected (iv) ensures that the obtained prediction intervals are informative in the sense that they do not include benchmark values (i.e., values in $[a,b]$). This ensures that the selection method is meaningful for the considered aim, which formally entails the FDR control (ii). Obviously, since InfoSCOP is an InfoSP method for preprocessed $p$-values (24), a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3.
Corollary 4.1 is illustrated on Figure 3 when $\mathcal{S}^{(0)}$ is taken here as $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ (with the score $S_{y}(x)=|\mu(x)-y|/\sigma(x)$). In the first row, errors appear more often near to the selection area. Hence, while the marginal prediction intervals control the FCR at level $\alpha$ when selecting all the covariates (as granted by (9)), the FCR is inflated for a naive selection that selects each example with a prediction interval at level $\alpha$ not intersecting $[a,b]$ (that is, naive selection is given by $\mathcal{S}_{1}$ in the recursion of Remark 3.1). To maintain the FCR control at level $\alpha=0.1$, InfoSP adjusts the width of the prediction interval in an accurate way to accommodate the informative constraint. InfoSCOP is roughly the same as InfoSP in this case, because the largest scores are kept in the calibration sample after initial selection. This is the most unfavorable setting for InfoSCOP, but it nevertheless performs similarly to InfoSP. In the second row, errors appear less often on the selection area, which makes the FCR of the naive selection and InfoSP too conservative. By contrast, the initial selection of InfoSCOP removes the largest scores of the calibration sample, resulting in much narrower prediction intervals, and thus in a much larger resolution-adjusted power (even better than the naive procedure).
Classical conformal InfoSP InfoSCOP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Informative prediction intervals when excluding $[a,b]$ (homoscedastic Gaussian regression model with perfect variance prediction), see text. First row: the predictor $\mu$ (dotted line) does not approximate well the true $\mu^{*}(x)=\mathbb{E}[Y|X=x]$ (solid line) in the selection area. Second row: the predictor $\mu$ (dotted line) does no approximate well the true $\mu^{*}$ (solid line) out of the selection area. The marginal prediction intervals are depicted in light blue, while the new informative prediction intervals (InfoSP and InfoSCOP) are depicted in red. While the plot corresponds to one data generation, the FCR and adjuster power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.
Remark 4.1.
In Corollary 4.1, we consider the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ for simplicity of exposition, but we can use any score function satisfying Assumption 1. For instance, for the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$, the corresponding $q_{i}$’s have the expression
$q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+\overline{p}^{(\mu(X_{n+i}))}_{i}{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}},$
where $\mu(x)=(q_{\beta_{0}}(x)+q_{\beta_{1}}(x))/2$ and where the $\overline{p}^{(y)}_{i}$’s are computed by using this score function. This leads to the prediction intervals $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (which do not contain $[a,b]$), by using the notation of Corollary 4.1.
4.2 Length-restricted prediction intervals
[…]
Corollary 4.2.
[…]
  • (i)
    InfoSP selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ with $q_{i}$ given by (12).
  • (ii)
    The selected prediction intervals are of length at most $2\lambda_{0}$ and are of the form $\mathcal{C}_{n+i}=[\mu(x)-S_{(n_{\alpha}(\mathbf{p}))}\sigma(x),\mu(x)+S_{(n_{\alpha}(\mathbf{p}))}\sigma(x)]$, where $S_{(1)}\leq\dots\leq S_{(n)}$ are the ordered calibration scores $S_{Y_{j}}(X_{j})$, $1\leq j\leq n$ (with $S_{(n+1)}=+\infty$), and $n_{\alpha}(\mathbf{p})=\lceil(1-\alpha|\mathcal{S}(\mathbf{p})|/m)(n+1)\rceil$.
  • (iii)
    These prediction intervals control the FCR at level $\alpha$ in the sense of (17).
A similar result holds for InfoSCOP. In Corollary 4.2 (ii), the length of the prediction interval on the selection is always granted to be (at most) of the correct size, even if adjusting the level is necessary to account for selection (which de facto enlarge the prediction interval). Thanks to the $\mathrm{BH}(\mathbf{q})$ selection the size-adjustment is automatic, while maintaining the FCR control.
Proof.
By Remark 2.1, we have $q_{i}\leq\alpha$ iff $\sigma(X_{n+i})S_{(\lceil(1-\alpha)(n+1)\rceil)}\leq\lambda_{0}$. This implies the expression of $q_{i}$ and that $|\mathcal{C}_{n+i}|=2\sigma(X_{n+i})S_{(\lceil(1-\alpha)(n_{\alpha}(\mathbf{p})+1)\rceil)}\leq 2\lambda_{0}$ since $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$. ∎
Figure 4 displays length-restricted informative prediction intervals in particular settings. In the first row, errors are more likely to occur on the selection (due to under-estimation of the variance), while in the second row, errors are less likely to occur on the selection (due to over-estimation of the variance). Hence, the comment is similar to the previous section: InfoSP and InfoSCOP are similar in the first situation but InfoSCOP improves InfoSP in the second.
Remark 4.2.
Corollary 4.2 easily extends to the case of conformalized quantile regression, by considering the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$. In that case, $q_{i}\leq\alpha$ iff $2S_{(\lceil(1-\alpha)(n+1)\rceil)}+q_{\beta_{1}}(X_{n+i})-q_{\beta_{0}}(X_{n+i})\leq 2\lambda_{0}$ which leads to the formula (13) for the $q_{i}$’s and to the prediction intervals $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (of length at most $2\lambda_{0}$), with the notation of Corollary 4.2.
Classical conformal InfoSP InfoSCOP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Informative prediction intervals when length-restricted (heteroscedastic Gaussian regression model with perfect mean prediction), see text. First row: the predictor $\sigma$ under-estimate the true $\sigma^{*}(x)=\mathbb{V}^{1/2}[Y|X=x]$ in the selection area. Second row: the predictor $\sigma$ over-estimate the true $\sigma^{*}$ in the selection area. The marginal prediction intervals are depicted in light blue, while the new informative prediction intervals (InfoSP and InfoSCOP) are depicted in red. While the plot corresponds to one data generation, the FCR and adjuster power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.
5 Application to classification
Here we consider the classification case $\mathcal{Y}=[K]$, apply our results and provide some new ones. We deal here both with the iid model and the class-conditional model.
5.1 Post-processing
In classification, any selective prediction set procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is post-processed by setting, for $i\in\mathcal{S}$, $\mathcal{C}_{n+i}=\arg\min_{k\in[K]}\{S_{k}(X_{n+i})\}$ whenever $\mathcal{C}_{n+i}=\emptyset$ (that is, if empty take the smallest non-conformity score). Clearly, this operation can only decrease the FCP while it can only increase the adjusted power, so it should always be preferred in the classification case. In this paper, InfoSP and InfoSCOP always refer to the post-processed procedures in the classification case.
5.2 Choosing the appropriate $p$-value collection in classification
While the family of full-calibrated $p$-values are only valid for the iid model, the family of class-calibrated $p$-values are valid both in the iid and in the class-conditional model.
Thus, for the iid model, we can in principle use either full-calibrated $p$-values or class-calibrated $p$-values. In the applications we consider next, it appears that using full-calibrated $p$-values in InfoSP is best. We support this claim by theory for non-trivial classification in § 5.4, and by numerical experiments in § 5.3, § 5.4, and § F. We note that for the initial selection step in InfoSCOP, class-calibrated $p$-values can be useful, as demonstrated in § 5.3.
For the class-conditional model, there can be a label shift from the calibration to the test sample. So the full-calibrated $p$-values are not valid, and class-calibrated $p$-values must be used. In § B we consider more generally weighted class-calibrated $p$-values, where the weights are functions of estimators of the proportion of labels in each class in the test sample.
5.3 An illustrative example: prediction sets excluding a null class
Suppose the analyst is interested in reporting prediction sets that exclude a null class, say class $y_{0}=1$ (see first item of Examples 1.2 and 2.2). We consider the following novel procedures, in addition to the naive procedure using the classic conformal procedure, that reports $\mathcal{C}^{\alpha}_{n+i}$ only if $\mathcal{C}^{\alpha}_{n+i}$ does not contain the null class:
  • InfoSP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$.
  • InfoSCOP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})$, with initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ being the BH procedure applied to the class-calibrated adaptive $p$-value family $(\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}},i\in[r+1,n+m])$, given by
    $\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}}=\hat{\pi}_{1}\tilde{p}^{(1)}_{i},\>i\in[r+1,n+m],$
    using $\{(X_{j},Y_{j}),j\in[r]\}$ and $\{(X_{j},Y_{j}),j\in[r+1,n+m]\}$ as calibration and test samples, respectively, and with calibrated-based estimator $\hat{\pi}_{1}=\frac{\sum_{i=1}^{r}{\mathbf{1}\{Y_{i}=1\}}+1}{r+1}$.
  • InfoSP on class-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}})$.
Note that for InfoSCOP above uses class-calibrated $p$-values for initial selection (because these are better to detect examples from the non-null class), and full-calibrated $p$-values on the selected examples from $[r+1,n+m]$ in the second step (because these are better $p$-values for building prediction sets in the iid model).
We consider a Gaussian mixture model with $K=3$ components, where each component is bivariate normal. The centers for components 1,2,3 are, respectively, (0,0), (SNR,0), (SNR,SNR). So the overlap between components is larger as SNR decreases. We consider the case of balanced classes in the calibration sample, as well as the case of unbalanced classes where the null class is much larger than the others. Specifically, the balanced case has class probabilities 0.33, 0.33, and 0.34, and the unbalanced case has class probabilities 0.15, 0.10, and 0.75 (the null class). In the balanced case, we consider the iid setting where the test sample has the same class probabilities as the calibration sample (depicted in Figure 5 for an SNR value of 3), as well as class conditional setting where the test sample has class probabilities 0.2,0.2, and 0.6 (the null class), so the label shift is large. We estimate the probability of being in each class with a support vector classifier implemented by the e1071 R package Meyer et al., (2023).
Figure 6 shows the FCR and resolution-adjusted power of all the procedures considered. As expected, the classic conformal procedure does not control the FCR for any data generating model (it uses the full-calibrated $p$-values in the iid setting, and the class-calibrated $p$-values in the class conditional setting). All other procedures control the FCR.
For the iid model, InfoSCOP on full calibrated $p$-values has better power than the alternatives for prediction sets excluding a null class. Its advantage over InfoSP is primarily due to the fact that after pre-processing, almost all test examples are non-null, as illustrated in Figure 5 for a single data generation. The differences between the procedures are qualitatively the same, but even greater, when the overlap between components is larger, see Figure 13 in the SM. This is because when the overlap with the null class is large, after pre-processing only examples with better scores are considered, as illustrated in Figure 14 in the SM. For completeness, we also provide InfoSP on class calibrated $p$-values in the iid setting, to demonstrate numerically the potentially large power advantage from using the full calibrated $p$-values over the class calibrated $p$-values.
For the class conditional model, InfoSP has lower power than classic conformal, but the power is reasonable. In Figure 9 in the SM we compare it to two procedures that weigh the classes according to their estimated relative frequencies.
Figure 5: The test sample in a single data generation for selecting informative prediction sets that exclude a null class. The setting is that of unbalanced classes, and the SNR is 3. The data points from the test sample of each of the three classes, where the null group is in blue: left panel for the entire test sample, right panel the remaining examples from the test set after the initial selection step.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Selecting prediction sets excluding a null class in a classification setting. FCR (top row), and resolution-adjusted power (bottom row) versus SNR. The iid setting in columns 1 and 2, with balanced classes and unbalanced classes, respectively. The class conditional setting in column 3, with a large label shift: the class probabilities are equal in the calibration sample and 0.2,0.2, and 0.6 (the null class) in the test sample. The number of data generations was 2000, 1000 data points were used for training, and $n=m=500$. See details of the data generation in § 5.3.
5.4 Selecting non-trivial prediction sets
Suppose the analyst is interested in reporting prediction sets that are not equal to $[K]$ (see first item of Examples 1.2 and 2.2). In that case, we argue that InfoSP has an FCR close to $\alpha$ in the iid model. Intuitively, this comes from the selection rule $\mathcal{S}=\mathrm{BH}(\mathbf{\overline{p}})$ which is such that
${\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}}),i\in\mathcal{S}\}}={\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}})\}}.$
It means that it is not possible to fail to cover at the adjusted level without being selected. This is not the case in other selection rules, e.g., excluding a null class, where it is possible that the true class label is not covered at the adjusted level even if the example is not selected (thus implying that the adjusted level is conservative, since for FCR control we guard against non-coverage at the adjusted level for all examples). We formalize fully the argument for $K=2$ in the following result.
Proposition 5.1.
In the iid classification model with $K=2$, consider the non-trivial informative subset collection =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq 1\}$ and assume that the score functions satisfy Assumption 1 with $\sum_{k\in[K]}S_{k}(x)=1$ and $S_{k}(x)\geq 0$. Let $p_{0}$ be the probability that $S_{Y_{i}}(X_{i})$ is the maximum score $\max\{S_{1}(X_{i}),S_{2}(X_{i})\}$. Then we have
$\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}),P_{X,Y})=\alpha(1-(1-p_{0})^{n+1}),$
provided that $(n+1)\alpha/m$ is an integer.
The proof is given in § C.4, which also shows that $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$ coincides with the procedure of Zhao and Su, (2023) for $K=2$. In typical applications (where classes are not very well separated) the value of $(1-(1-p_{0})^{n+1})$ is close to one, which means that the FCR of our procedure should be close to $\alpha$, at least for $K=2$.
To complement our theoretical result, we provide numerical results in Figure 7. Clearly, InfoSCOP is not improving InfoSP for non-trivial selection. InfoSP has the best power, with InfoSCOP a close second, on full calibrated $p$-values. InfoSP on class calibrated $p$-values has much lower power in the unbalanced setting. As expected, the classic conformal procedure does not control the FCR and the level for InfoPS is about $0.05$ for a range of SNR values. All other procedures control the FCR.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Selecting non-trivial prediction sets in the classification iid model, in the case of balanced classes (top row) and unbalanced classes (bottom row), for the Gaussian mixture model with $K=3$ classes described in § 5.3. FCR (first column), resolution-adjusted power (second column) and the expected fraction of covering prediction sets (third column) versus SNR. The number of data generations was 2000, 1000 data points were used for training, and $n=m=500$.
6 Application to directional FDR control
Consider for this section that we have $(X_{i},Z_{i})_{i\in[n+m]}$ with real-valued outcomes $Z_{i}\in\mathbb{R}$, and we aim at deciding if $Z_{n+i}\in[a,b]$, $Z_{n+i}<a$ or $Z_{n+i}>b$ (without producing prediction intervals for the $Z_{n+i}$), for two benchmark values $a\leq b$. More formally, we want to build a selection $\mathcal{S}\subset[m]$ and a (point-wise) decision $\widehat{Y}\in\{1,2,3\}$ from the observed samples such that
$\mathrm{FDR}_{{\tiny\mbox{dir}}}(\mathcal{S},\widehat{Y}):=\sup_{P_{X|Y},Y}\mathbb{E}_{X\sim P_{X|Y}}\left[\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\neq\widehat{Y}_{n+i}\}}}{1\vee|\mathcal{S}|}\right]\leq\alpha,$ (26)
where $Y_{j}={\mathbf{1}\{Z_{j}<a\}}+2{\mathbf{1}\{Z_{j}\in[a,b]\}}+3{\mathbf{1}\{Z_{j}>b\}},$ $j\in[n+m]$. Our theory yields the following result.
Corollary 6.1.
Consider the class-conditional (classification) model on $(X_{i},Y_{i})_{i\in[n+m]}$ with $K=3$ classes. Consider any score function satisfying Assumption 1 (in this classification model). Consider the procedure that selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ with
$q_{i}=\max(\tilde{p}^{(2)}_{i},\min(\tilde{p}^{(1)}_{i},\tilde{p}^{(3)}_{i})),\>i\in[m],$
where $\tilde{p}^{(y)}_{i}$ are the class-conditional $p$-values computed as in (8), and with the decision
$\widehat{Y}_{n+i}$ $={\mathbf{1}\{S_{1}(X_{n+i})\geq S_{3}(X_{n+i})\}}+3{\mathbf{1}\{S_{3}(X_{n+i})>S_{1}(X_{n+i})\}},i\in\mathcal{S}.$
Then this procedure controls the directional FDR at level $\alpha$ in the sense of (26).
Proof.
We consider =$\mathcal{I}=\{C\subset[K]\>:\>y_{0}\notin C,|C|\leq 1\}$ for $y_{0}=2$ (see Example 2.3) in the classification setting based on the sample $(X_{j},Y_{j})$’s ($K=3$), and we note that the FCR coincides with the directional FDR in that case. Hence, the result comes directly from Theorem 3.1 (note that InfoSP is post-processed here, see § 5.1). ∎
Remark 6.1.
For $b=a$ and continuous outcomes (with, say, $\tilde{p}^{(2)}_{i}=0$), there are almost surely only two classes, $Z_{i}=1$ corresponding to $Y_{i}<a$ and $Z_{i}=3$ corresponding to $Y_{i}>a$. The procedure is thus InfoSP for non-trivial classification, with class-calibrated $p$-values for $K=2$. This procedure coincides with the directional FDR procedure in Guo and Romano, (2015) applied to conformal $p$-values. The proof of validity in Guo and Romano, (2015) assumes that the test statistics for the $m$ hypotheses are independent. Interestingly, with the dependence induced by the conformal $p$-values, the same procedure is still valid, as formalized in Corollary 6.1.
7 Informative prediction sets for 3 classes of animals
In this section, we illustrate the performance of our methods on real data. We use the image dataset CIFAR-10 (https://www.cs.toronto.edu/~kriz/cifar.html), which consists of 60000 32x32 colour images in 10 classes, with 6000 images per class. We restrict the analysis to 3 classes: birds, cats and dogs.
We consider 4 scenarios: a) non-trivial classification task in a i.i.d. setting, b) non-trivial classification task in a class-conditional setting, c) non-null classification task in a i.i.d. setting, and d) non-null classification task in a class-conditional setting (in c) and d), the null class is the bird class). Specifically, while the classes are in equal proportions by default, we introduce a label shift between the calibration sample and the test sample for the class-conditional settings, in b) by modifying the classes proportions of the calibration sample to 20% for the bird class versus 40% for the cat/dog class and in d) by modifying the classes proportions of the test sample to 50% for the bird class versus 25% for the cat/dog class. In each experiment, the test size is $m=5000$ and the calibration size is $n=5000$.
In the i.i.d. settings (scenarios a) and c)), we evaluate the procedures InfoSP and InfoSCOP with full-calibrated $p$-values and in the class-conditional settings (scenarios b) and d)) we evaluate InfoSP with class-calibrated $p$-values, each being also compared with classical conformal prediction (denoted by CC). For all procedures the non-conformity score is $S_{y}(x)=1-\pi_{y}(x)$ where $\pi_{y}(x)$ is an estimator of the probability that the class of $x$ is $y$ and is learned using a convolutional neural network (CNN) with 2 convolutional layers, one pooling layer, and 3 fully-connected layers, trained for 20 epochs with a learning rate of 0.01 on a sample size of 5000. To assess the power of the methods, in addition to the resolution-adjusted power, we plot the selection rate (SR) of the procedures, defined as the average proportion of informative prediction sets returned, and the average length of the informative prediction sets.
In each setting, the FDR and each power metric for the methods are evaluated by using 100 runs and the results are reported in Figure 8 for $\alpha=0.1$. The conclusions are qualitatively similar to the experiments of § 5: in all settings, classical conformal prediction yields a FCR that severely exceeds the marginal level with an inflation of about 50%. By contrast, our procedures InfoSP and InfoSCOP control the FCR at the target nominal level, both without and with label shift. In terms of power, concerning the i.i.d. settings, pre-processing is not useful for non-trivial classification as expected and the performances of InfoSP and InfoSCOP are similar with a FCR close to the nominal level for both in that case. Conversely, for the non-null classification task, InfoSP is conservative while InfoSCOP is more powerful and displays a FCR close to $\alpha$. When there is label shift, in the case of non-trivial classification InfoSP displays a FCR close to $\alpha$. In the non-null classification case, however, the label shift increases the difficulty of the task in the sense that for a fixed selection, under-covering the null class in the test sample results in more false informative sets. Hence, the power is low in that case. Finally, in all scenarios, while our procedures InfoSP and InfoSCOP output a lower number of informative sets compared to CC, which is necessary in order to control the FCR, the average size of the prediction sets that are informative is comparable.
a) Non-trivial classification
b) Non-trivial classification with label shift
c) Non-null classification
d) Non-null classification with label shift
Figure 8: FCR, SR, average length on selection and resolution-adjusted power for the methods, for $\alpha=0.1$.
8 Conclusion and discussion
In this paper, we have introduced new methods for providing conformal prediction sets after selection with controlled FCR, that impose a user-specific constraint on the produced prediction sets, corresponding to a collection of so-called informative subsets $\mathcal{I}$. In contrast with previous literature in the field, the selection and prediction stages are intertwined, which results in a BH-type selection procedure on adjusted $p$-values (the $q_{i}$’s) that can further be explicitly derived in specific settings and that by definition produces prediction sets satisfying the desired constraint of belonging to $\mathcal{I}$.
Our methods are very general, and they are relevant to applications in classification and regression. We showed examples in § 4-§ 7 for informative subsets of interest. We used common scores for the examples, but many other scores from the literature can be used with our suggested procedures InfoSP and InfoSCOP as long as Assumption 1 is satisfied. In addition, transfer learning scores can also be handled by our theory (see Assumption 6 in SM). This is known to greatly improve the conformal inference when there is a domain shift between the learning sample and the calibration+test samples (Courty et al.,, 2017; Gazin et al.,, 2023) and developing specific application cases for the latter is of interest for future investigations.
For the iid model, InfoSCOP improves over InfoSP in all considered examples, except when selecting non-trivial prediction subsets in classification (Example 1.2 item 2), for which we establish that InfoSP almost exhausts the FCR level (see Proposition 5.1for $K=2$). The procedure InfoSCOP splits the calibration sample in order to apply an efficient initial selection step on part of the calibration sample and the test sample. When the calibration sample is small, it may be of interest to examine whether alternative methods (e.g., relevant variants of the Jackknife Barber et al.,, 2021) may be incorporated into our framework in settings where InfoSCOP improves over InfoSP but splitting may be too costly. When the calibration sample is large enough, as in our numerical examples, splitting the calibration sample for the initial selection $S^{(0)}$ worked well. However, there are many possibilities about how to perform the initial selection. The choice is important because it defines the pre-processed $p$-values (24) that can be seen as $p$-values “conditionally on being selected”. Different ways of choosing $S^{(0)}$ have been investigated: trying to rule out all the examples in null class (§ 5.3) or trying to mimic the $\mathrm{BH}(\mathbf{q})$ selection that will be applied at the second stage to reduce the selection effect (§ 4). Finding an optimal way of calibrating $S^{(0)}$ is an interesting avenue for future research.
For non-trivial prediction subsets in classification, InfoSP is optimal when $K=2$ for oracle scores, i.e., $S_{k}(X_{j})=\mathbb{P}(Y_{j}\neq k\mid X_{j}),k\in[K],j\in[n+m]$. Specifically, Zhao and Su, (2023) showed that their classification procedure, which coincides with InfoSP for $K=2$, is optimal for controlling the expected number of non-covering prediction sets divided by the expected number of selected examples, denoted by mFSR in their paper. An open question is whether InfoSP for $K>2$ is optimal when the scores are oracle scores for the resolution adjusted power objective or a variant thereof. More generally, developing an optimality theory for selective informative prediction sets (for non-trivial prediction sets as well as for other notions informativeness) is of great interest.
We provided a class conditional variant of InfoSP, with class-conditional guarantees. We proved that our strategy can be followed in the case where the classes of the calibration and test samples are arbitrary fixed, even when the class proportions in the calibration are very different than in the test. The main point is that working with the class-calibrated $p$-value collection allows to maintain the FCR control in this strong sense. In § B we suggest additionally weighted procedures, that incorporate the estimated class proportions. These procedures are not necessarily more powerful than InfoSP, and further research is needed in order to make recommendations about when to use the weighted procedures, and about weight adaptation to the specifics of the data.
Acknowledgments
The authors acknowledge the grants ANR-21-CE23-0035 (ASCAI) and ANR-23-CE40-0018-01 (BACKUP) of the French National Research Agency ANR, and the Emergence project MARS of Sorbonne Université. The second author was supported in part by Israeli Science Foundation grant no. 2180/20.
Appendix A FCR control for BY-type selection rules
In this section, we present a general approach for FCR control (§ A.1), which relies on the following:
  • a general class of $p$-values, including both iid and class-conditional conformal $p$-values (§ A.3);
  • a general selection rule, including informative selection rules (§ A.4).
The assumption on the selection rule is inspired by Benjamini and Yekutieli, (2005); Benjamini and Bogomolov, (2013).
A.1 General statement
Let $m\geq 1$, $\mathcal{Y}\subset\mathbb{R}$ and consider a family $\mathbf{p}=\{p^{(y)}_{i},i\in[m],y\in\mathcal{Y}\}$ of random variables taking values in $[0,1]$ and a vector $Y=(Y_{n+i},i\in[m])$ taking values in $\mathcal{Y}$. We make use of the notation $\mathbf{p}_{-i}:=(p^{(y)}_{j})_{j\neq i,y\in\mathcal{Y}}$ for all $i\in[m]$.
First, we introduce the following assumption on $\mathbf{p}$ and $Y$:
Assumption 4.
There exists a vector $W=(W_{i},i\in[m])$ of multivariate random variables with
  • (i)
    for all $i\in[m]$, the random vector $\mathbf{p}_{-i}=(p^{(y)}_{j})_{j\neq i,y\in\mathcal{Y}}$ can be almost surely written as $\Psi_{i}(p^{(Y_{n+i})}_{i},W_{i})$ where $u\in[0,1]\mapsto\Psi_{i}(u,W_{i})\in\mathbb{R}^{[m-1]\times\mathcal{Y}}$ is a nondecreasing function (in a coordinate-wise sense for the image space).
  • (ii)
    for all $i\in[m]$, the following super-uniformity property holds
    $\mathbb{P}(p_{i}^{(Y_{n+i})}\leq t\>|\>W_{i})\leq t,\>\>t\in[0,1].$ (27)
Second, for any selection rule $\mathcal{S}\subset[m]$, that is, any measurable function of $\mathbf{p}$ valued in the subsets of $[m]$, we introduce the quantity (similar to Benjamini and Yekutieli, (2005); Bao et al., (2024))
$s_{i}^{\min}(\mathbf{p}_{-i})=\min_{z\in[0,1]^{\mathcal{Y}}:i\in\mathcal{S}(z,\mathbf{p}_{-i})}\left|\mathcal{S}(z,\mathbf{p}_{-i})\right|\geq 1,\>\>\>i\in[m],$ (28)
with the following assumption.
Assumption 5.
For $i\in[m]$, $s_{i}^{\min}(\mathbf{p}_{-i})$ is a coordinate-wise nonincreasing function of $\mathbf{p}_{-i}$.
Theorem A.1.
Let us consider a $p$-value family $\mathbf{p}$ and a label/outcome vector $Y$ satisfying Assumption 4, for a selection rule $\mathcal{S}\subset[m]$ with $s_{i}^{\min}=s_{i}^{\min}(\mathbf{p}_{-i})$ (28) satisfying Assumption 5. Then, the procedure $\mathcal{R}_{\alpha}=\big{(}\mathcal{C}^{\alpha s_{i}^{\min}/m}_{n+i}(\mathbf{p})\big{)}_{i\in\mathcal{S}}$, for which the prediction set is defined as in (4) with level $\alpha s_{i}^{\min}/m$, satisfies $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha,$ for which the expectation $\mathbb{E}$ is taken w.r.t. the same probability than the one of (27) in Assumption 4 (ii).
The inequality $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha$ in Theorem A.1 will be used both in the cases where $Y_{n+i}$ is fixed (conditional model) or not (iid model). The proof is provided in § A.2. Note that the considered assumptions make the proof particularly simple. Assumptions 4 and 5 are studied in § A.3 and § A.4, respectively.
A.2 Proof of Theorem A.1
By definition (2), we have
$\mathrm{FCP}(\mathcal{R}_{\alpha},Y)$ $=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}_{n+i}^{\alpha s_{i}^{\min}/m}(\mathbf{p})\}}}{1\vee|\mathcal{S}|}=\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{1\vee|\mathcal{S}|}$
$\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}}\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}},$
where the first inequality follows from definition (28) and the second inequality follows from ignoring the fact that $i\in\mathcal{S}$. This entails
$\mathbb{E}\left(\mathrm{FCP}(\mathcal{R}_{\alpha},Y)\right)$ $\leq\sum_{i\in[m]}\mathbb{E}\left(\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}}\right)$
$=\sum_{i\in[m]}\mathbb{E}\left(\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}(\mathbf{p}_{-i})/m\}}}{s^{\min}_{i}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\right),$
by using the random vector ${W}=(W_{i},i\in[m])$ defined in Assumption 4. Now, combining Assumption 4 (i) with Assumption 5, we have that
$s^{\min}_{i}(\mathbf{p}_{-i})=s^{\min}_{i}(\Psi_{i}(p^{(Y_{n+i})}_{i},W_{i}))$
is a nonincreasing function of $p^{(Y_{n+i})}_{i}$. By Assumption 4 (ii) and applying Lemma D.1 (conditionally on $W_{i}$), we obtain
$\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}(\mathbf{p}_{-i})/m\}}}{s^{\min}_{i}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\leq\frac{\alpha}{m},$
Putting this back into the FCR bound implies the result.
A.3 Examining Assumption 4
We show here that the $p$-value families we consider satisfy Assumption 4 both in the iid and conditional model. For this, it is interesting to relax Assumption 1 by assuming that the score functions can use the covariates of the calibration+test samples in an exchangeable way, as suggested in Marandon et al., (2024); Gazin et al., (2023). This is useful for instance when the learning sample and the calibration+test samples are not based on the same distribution, so that the scores may be improved by using transfer learning; we refer to Gazin et al., (2023) for more details on this.
Assumption 6.
[…]
$p$-value family $\mathbf{\overline{p}}$
The $p$-value family $\mathbf{\overline{p}}$ given by (7) satisfies Assumption 4 in the iid model.
Proposition A.2.
Let us consider a model where the variables $(X_{i},Y_{i})$, $i\in[n+m]$, are exchangeable and score functions satisfying Assumptions 6. Then the $p$-value family $\mathbf{\overline{p}}$ given by (7) satisfies Assumption 4.
Proof.
Let us first establish Assumption 4 (i) by following an argument similar to the one of Bates et al., (2021); Marandon et al., (2024). By Assumption 6, we can work on an event where the elements of the sets $A_{i}=\{S_{Y_{k}}(X_{k}),k\in[n]\}\cup\{S_{Y_{n+i}}(X_{n+i})\}$ are all distinct. For any $i\in[m]$, we have by (7) that for all $j\in[m]\backslash\{i\}$ and $y\in\mathcal{Y}$,
$\overline{p}^{(y)}_{j}$ $=\frac{1}{n+1}\Big{(}1+\sum_{k=1}^{n}{\mathbf{1}\{S_{Y_{k}}(X_{k})\geq S_{y}(X_{n+j})\}}\Big{)}$
$=\frac{1}{n+1}\Big{(}{\mathbf{1}\{S_{Y_{n+i}}(X_{n+i})<S_{y}(X_{n+j})\}}+\sum_{s\in A_{i}}{\mathbf{1}\{s\geq S_{y}(X_{n+j})\}}\Big{)}.$
Denoting $A_{i}=\{a_{i,(1)},\dots,a_{i,(n+1)}\}$ with $a_{i,(1)}>\dots>a_{i,(n+1)}$, and noting that $S_{Y_{n+i}}(X_{n+i})=a_{i,(\ell)}$ with $\overline{p}^{(y)}_{i}=\ell/(n+1)$, we may write
$\mathbf{\overline{p}}_{-i}:=(\overline{p}^{(y)}_{j})_{j\in[m]\backslash\{i\},y\in\mathcal{Y}}=\Psi(\overline{p}^{(Y_{n+i})}_{i},\overline{W}_{i}),$
by letting for $u\in(0,1]$,
$\overline{W}_{i}$ $:=(A_{i},(S_{y}(X_{n+j}))_{j\in[m]\backslash\{i\},y\in\mathcal{Y}})$
$\Psi(u,\overline{W}_{i})$ $:=\left(\frac{1}{n+1}\Big{(}{\mathbf{1}\{a_{i,(\lceil u(n+1)\rceil)}<S_{y}(X_{n+j})\}}+\sum_{s\in A_{i}}{\mathbf{1}\{s\geq S_{y}(X_{n+j})\}}\Big{)}\right)_{j\in[m]\backslash\{i\},y\in\mathcal{Y}}.$
Clearly, each of the elements inside $\Psi(u,\overline{W}_{i})$ is nondecreasing in $u$, which gives Assumption 4 (i) (note that $\Psi$ does not depend on $i$ in this context).
Next, we establish Assumption 4 (ii). Since the variables $(X_{i},Y_{i})$, $i\in[n+m]$, are exchangeable and since the scores functions, the set $A_{i}$, and $(S_{y}(X_{n+j}))_{j\in[m]\backslash\{i\},y\in\mathcal{Y}}$ are invariant by permutations of $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X_{n+i},Y_{n+i})$ (by using Assumptions 6), we have that the random vector $(S_{Y_{1}}(X_{1}),\dots,S_{Y_{n}}(X_{n}),S_{Y_{n+i}}(X_{n+i}))$ is exchangeable conditionally on $\overline{W}_{i}$. Since there are no ties in the vector, it follows that $(n+1)\overline{p}^{(Y_{n+i})}_{i}$ (i.e., the rank of $S_{Y_{n+i}}(X_{n+i})$ in $A_{i}$) is uniformly distributed in $[n+1]$ conditionally on $\overline{W}_{i}$. Thus Assumption 4 (ii) is satisfied (note that the conditional probability is well defined thanks to the regularity condition in Assumption 6). ∎
$p$-value family $\mathbf{{\widetilde{p}}}$
The $p$-value family $\mathbf{{\widetilde{p}}}$ given by (8) satisfies Assumption 4 in the conditional model.
Proposition A.3.
In the case $\mathcal{Y}=[K]$, let us consider score functions satisfying Assumptions 6 and a model for the variables $(X_{i},Y_{i})$, $i\in[n+m]$, for which $(Y_{i},i\in[n+m])$ is a deterministic vector and, for each $y\in\mathcal{Y}$, the variables $(X_{i})_{i\in[n+m]:Y_{i}=y}$, are exchangeable. Then the $p$-value family $\mathbf{{\widetilde{p}}}$ given by (8) satisfies Assumption 4.
The proof is similar to the proof of Proposition A.2. We provide it below for completeness.
Proof.
For any $i\in[m]$, we have by (8) that for all $j\in[m]\backslash\{i\}$ and $y\in\mathcal{Y}$,
$\tilde{p}^{(y)}_{j}$ $=\frac{1}{|\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}|+1}\Big{(}1+\sum_{k\in\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}}{\mathbf{1}\{S_{y}(X_{k})\geq S_{y}(X_{n+j})\}}\Big{)}$
$=\frac{1}{n^{(y)}_{i}}\Big{(}{\mathbf{1}\{S_{y}(X_{n+i})<S_{y}(X_{n+j})\}}+\sum_{s\in A^{(y)}_{i}}{\mathbf{1}\{s\geq S_{y}(X_{n+j})\}}\Big{)},$
by letting $n^{(y)}_{i}=|A^{(y)}_{i}|$ and $A^{(y)}_{i}=\{S_{y}(X_{k}),k\in\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}\}\cup\{S_{y}(X_{n+i})\}$ whose elements can be assumed to be all distinct by Assumptions 6 (strictly, this is only true for labels $y\in\mathcal{Y}$ that appears at least once in the fixed sample $(Y_{i},i\in[n+m])$, but the labels $y$ not appearing in $(Y_{i},i\in[n+m])$ can be trivially handled because they correspond to $p$-values all equal to $1$). Denoting $A^{(y)}_{i}=\{a^{(y)}_{i,(1)},\dots,a^{(y)}_{i,(n^{(y)}_{i})}\}$ with $a^{(y)}_{i,(1)}>\dots>a^{(y)}_{i,(n^{(y)}_{i})}$, and noting that $S_{y}(X_{n+i})=a^{(y)}_{i,(n^{(y)}_{i}\tilde{p}^{(y)}_{i})}$, we may write
$\mathbf{{\widetilde{p}}}_{-i}:=(\tilde{p}^{(y)}_{j})_{j\in[m]\backslash\{i\},y\in\mathcal{Y}}=\Psi_{i}(\tilde{p}^{(Y_{n+i})}_{i},{\widetilde{W}}_{i}),$
by letting
${\widetilde{W}}_{i}:=\big{(}(A^{(Y_{n+i})}_{i},(S_{Y_{n+i}}(X_{n+j}))_{j\in[m]\backslash\{i\}});(X_{j},Y_{j})_{j\in[n+m]:Y_{j}\neq Y_{n+i}}\big{)}$ (29)
and for $u\in(0,1]$, $\Psi_{i}(u,{\widetilde{W}}_{i}):=(\Psi^{(y)}_{i}(u,{\widetilde{W}}_{i}))_{y\in\mathcal{Y}}$ where
$\Psi^{(y)}_{i}(u,{\widetilde{W}}_{i})$
:=$:=\left\{\begin{array}[]{cc}\Big{(}\frac{1}{n^{(y)}_{i}}\Big{(}{\mathbf{1}\{a_{i,(\lceil un^{(y)}_{i}\rceil)}<S_{y}(X_{n+j})\}}+\sum_{s\in A^{(y)}_{i}}{\mathbf{1}\{s\geq S_{y}(X_{n+j})\}}\Big{)}\Big{)}_{j\in[m]\backslash\{i\}}&\mbox{ if $y=Y_{n+i}$;}\\ (\tilde{p}^{(y)}_{j})_{j\in[m]\backslash\{i\}}&\mbox{ if $y\neq Y_{n+i}$.}\end{array}\right.$
Clearly, the elements inside $\Psi_{i}(u,{\widetilde{W}}_{i})$ are nondecreasing in $u$ which gives Assumption 4 (i).
Next, to establish Assumption 4 (ii), we use that by assumption the vector
$(S_{Y_{n+i}}(X_{n+i}),S_{Y_{n+i}}(X_{k}),k\in\mathcal{D}^{(Y_{n+i})}_{{\tiny\mbox{cal}}})$
is exchangeable conditionally on ${\widetilde{W}}_{i}$ (by permutation invariance of ${\widetilde{W}}_{i}$, which also comes from Assumptions 6). Since there are no ties in the vector, it follows that $n^{(Y_{n+i})}_{i}\tilde{p}^{(Y_{n+i})}_{i}$ (i.e., the rank of $S_{Y_{n+i}}(X_{n+i})$ in $A^{(Y_{n+i})}_{i}$) is uniformly distributed in $[n^{(Y_{n+i})}_{i}]$ conditionally on ${\widetilde{W}}_{i}$. Thus Assumption 4 (ii) is satisfied. ∎
A.4 Examining Assumption 5
Proposition A.4.
Assumption 5 holds for the informative selection rule $\mathcal{S}(\mathbf{p})=\mathrm{BH}(\mathbf{q})$ with $s_{i}^{\min}(\mathbf{p}_{-i})=|\mathcal{S}(\mathbf{p})|$ whenever $i\in\mathcal{S}(\mathbf{p})$.
Note that the above result is also true for selection rule of the type $\mathcal{S}(\mathbf{p})=\{i\in[m]\>:\>q_{i}\leq\tau\}$ for some fixed threshold $\tau$.
Proof.
First, a classical property of BH procedure is the following leave-one-out property: for all $i\in[m]$, $i\in\mathrm{BH}(\mathbf{q})$ if and only if $\mathrm{BH}(\mathbf{q})=\mathrm{BH}(\mathbf{q}^{0,i})$ where $\mathbf{q}^{0,i}$ is the vector $\mathbf{q}$ where the $i$-th coordinate has been replaced by 00, see for instance Ferreira and Zwinderman, (2006); Sarkar, (2008); Roquain and Villers, (2011); Ramdas et al., (2019). This implies
$s_{i}^{\min}(\mathbf{p}_{-i})$ $=\min_{z\in[0,1]^{\mathcal{Y}}:i\in\mathcal{S}(z,\mathbf{p}_{-i})}\left|\mathcal{S}(z,\mathbf{p}_{-i})\right|=|\mathrm{BH}(\mathbf{q}^{0,i})|.$
Hence, the result is proved as soon as $\mathbf{q}$ is proved to be coordinate-wise nondecreasing in each $p$-value trajectory. This holds by Lemma A.5. ∎
Lemma A.5.
Suppose Assumption 2, then, almost surely, $\mathbf{q}$ defined by (11) is a nondecreasing function of each $p$-value ${p}^{(y)}_{i}$, $i\in[m]$ and $y\in\mathcal{Y}$.
Proof.
Let $\mathbf{p}$ and $\mathbf{p}^{\prime}$ be two $p$-value collections with ${p}^{(y)}_{i}\leq{p}^{\prime(y)}_{i}$, for all $i\in[m]$ and $y\in\mathcal{Y}$, with corresponding values $\mathbf{q}$ and $\mathbf{q}^{\prime}$. Let $i\in[m]$. By definition, $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p}^{\prime})\in\mathcal{I}$ and $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p})\subset\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p}^{\prime})$. By Assumption 2 (i) (iii), we have $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p})\in\mathcal{I}$, which in turn implies $q_{i}\leq q^{\prime}_{i}$ by definition of $q_{i}$. ∎
Appendix B Procedures using weighted class-calibrated $p$-values
Our procedures do not take into account the proportion of labels in each class in the test sample. However, in the classification case, these proportions are estimable from the data, and the estimates can aid inference.
Namely, let $\pi_{k}=\sum_{i=1}^{m}{\mathbf{1}\{Y_{n+i}=k\}}/m$ the true proportion of label $k$ in the test sample, $k\in[K]$, and consider the following possible estimates:
  • Calibration-based estimator: $\hat{\pi}^{{\tiny\mbox{cal}}}_{k}=(|\mathcal{D}^{(k)}_{{\tiny\mbox{cal}}}|+1)/(n+1)=(1+\sum_{j\in[n]}{\mathbf{1}\{Y_{j}=k\}})/(n+1),$ $k\in[K]$;
  • Storey-$\lambda$ estimator: $\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}=\Big{(}1+\sum_{i=1}^{m}{\mathbf{1}\{p^{(k)}_{i}>\lambda\}}\Big{)}/(m(1-\lambda))$, $k\in[K]$.
The second one is similar to the classical estimator of true null hypotheses proportion in multiple testing (Storey,, 2002). Given the class-calibrated $p$-value family $\mathbf{{\widetilde{p}}}$ and one of the estimators $\hat{\pi}_{k}\in\{\hat{\pi}^{{\tiny\mbox{cal}}}_{k},\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}\}$, we define the corresponding adaptive (weighted) $p$-value collection $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt}}}=(\tilde{p}^{(k)}_{i,{\tiny\mbox{adapt}}},k\in[K],i\in[m]\}$ by
$\tilde{p}^{(k)}_{i,{\tiny\mbox{adapt}}}=\frac{\hat{\pi}_{k}}{w_{k}}\tilde{p}^{(k)}_{i},\>\>k\in[K],i\in[m],$ (30)
where $(w_{k},k\in[K])$ are deterministic nonnegative weights such that $\sum_{k\in[K]}w_{k}=1$. The rationale behind (30) is that the term $\hat{\pi}_{k}$ balances the false coverage errors between classes by trying to decrease $p$-values related to labels which do not appear much in the test sample. The weights $w_{k}$ are additional parameter that add flexibility, but they have to sum to one. If we use equal weights than the class-calibrated $p$-value is multiplied by $K\times\hat{\pi}_{k}$ which will be less than one only if $\hat{\pi}_{k}<1/K$.
Applying InfoSP with these adaptive $p$-values gives rise to a new procedure $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt}}})$ that we denote by Adapt-InfoSP.
Proposition B.1.
[…]
  • (i)
    in the class-conditional model,
    $\sup_{P_{X\mid Y},Y}\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}),P_{X\mid Y},Y)\leq\alpha\>\sum_{y\in[K]}w_{y}\frac{n+1}{m}\frac{\sum_{i\in[m]}{\mathbf{1}\{Y_{n+i}=y\}}}{\sum_{j\in[n]}{\mathbf{1}\{Y_{j}=y\}}+1}.$ (31)
  • (ii)
    in the iid model, $\sup_{P_{X,Y}}\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}),P_{X,Y})\leq\alpha$, that is, Adapt-InfoSP satisfies the FCR control (17).
Proposition B.1 is proved in § C.3. The bound (31) is only sharp when the labels are generated in the same way in the calibration and test sample (which implies the correct control in (ii)), so Adapt-InfoSP should not be used if the label proportions are expected to be (very) different between calibration and test samples.
We illustrate in Figure 9 the performance of the adaptive procedures for nonnull selection and non-trivial selection, respectively, in the set-up of unbalanced classes described in § 5.3. For the adaptive versions, $w_{k}=1/K$ for all $k\in[K]$, and $\lambda=1/2$. We consider two settings for the class conditional model: the test sample has class probabilities 0.1,0.1, 0.8 (i.e., a small label shift), and class probabilities 0.4,0.4,0.2 (i.e., a large label shift).
The only procedure with a theoretical class-conditional FCR guarantee is InfoSP on class calibrated $p$-values. The adaptive procedure with $\hat{\pi}_{k}^{cal},k\in[K]$ violates FCR control only when the label-shift is large for non-trivial selection. Interestingly, this procedure has excellent power when the label shift is small. The adaptive procedure with $\hat{\pi}_{k}^{Storey},k\in[K]$ is a close second in this case, but when the label shift is large it is no better than InfoSP in the settings considered. The fact that the adaptive procedure with $\hat{\pi}_{k}^{cal},k\in[K]$ tends to control the FCR (or inflate it only by a little), suggests (arguably) that for power purposes it may be reasonable to use it if the label shift is small.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: Selecting informative prediction sets in the class-conditional setting. FCR (left column), resolution-adjusted power (middle column), and the expected fraction of covering prediction sets (right column) versus SNR in a classification setting where prediction sets excluding a null class are of interest (top two rows) and when prediction sets excluding a trivial class are of interest (bottom two rows). The class probabilities in the calibration sample are 0.15, 0.1, and 0.75; in the test sample, we have a small label shift (rows 1 and 3) and a large label shift (rows 2 and 4). The number of data generations was 2000, and $n=m=500$.
Appendix C Proofs
C.1 Proof of Theorem 3.1
The proof is straightforward from the theory developed in § A. Namely, we apply Theorem A.1 by checking the two required assumptions: Assumption 4 holds for the two considered $p$-value collections (§ A.3); Assumption 5 holds for the considered $\mathrm{BH}(\mathbf{q})$ selection (§ A.4).
C.2 Proof of Theorem 3.3
The FCR control is a consequence of Theorem A.1: we only have to check that the pre-processed $p$-value family $\mathbf{\overline{p}}^{0}=(\overline{p}^{(0,y)}_{i},i+n\in\mathcal{S}^{(0)}\cap[n+1,n+m],y\in\mathcal{Y})$ (based on the conformal selection rule $\mathcal{S}^{(0)}$) satisfies Assumption 4 conditionally on $\mathcal{S}^{(0)}$. By Proposition A.2, it is sufficient to check that the variables $(X_{i},Y_{i})$, $i\in\mathcal{S}^{(0)}$, are exchangeable conditionally on $\mathcal{S}^{(0)}$. This is a direct consequence of the fact that the $(X_{j},Y_{j})$’s are iid and of the assumed permutation preserving property of $\mathcal{S}^{(0)}$.
Finally, the FDR control is a consequence of the FCR control by applying Lemma 2.1.
C.3 Proof of Proposition B.1
Let us first prove (i) by considering the class-conditional model. We follow the proof of Theorem A.1 (see § A.2) and we use that the $p$-value collection $\mathbf{{\widetilde{p}}}$ satisfies Assumption 4 (see Proposition A.3), where the probability in the super-uniform property (27) holds in the class-conditional model with ${\widetilde{W}}_{i}$, $i\in[m]$ given by (29). We also use that the informative selection rule $\mathcal{S}(\cdot)$ satisfies Assumption 5 (see Proposition A.4). Hence, following the same approach as in § A.2, and denoting $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}:=\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}$ and $\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}}:=(\tilde{p}^{(y)}_{j,{\tiny\mbox{cal}}})_{j\neq i,y\in\mathcal{Y}}$, we obtain
$\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}),P_{X\mid Y},Y)$ $\leq\sum_{i\in[m]}\mathbb{E}_{P_{X\mid Y}}\left(\mathbb{E}\left[\frac{{\mathbf{1}\{\tilde{p}^{(Y_{n+i})}_{i,{\tiny\mbox{cal}}}\leq\alpha s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})/m\}}}{s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})}\>\middle|\>{\widetilde{W}}_{i}\right]\right)$
$=\sum_{i\in[m]}\mathbb{E}_{P_{X\mid Y}}\left(\mathbb{E}\left[\frac{{\mathbf{1}\{\tilde{p}^{(Y_{n+i})}_{i}\leq\alpha s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})w_{Y_{n+i}}/(m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}})\}}}{s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})}\>\middle|\>{\widetilde{W}}_{i}\right]\right).$
Using now the super-uniform property (27), the fact that $\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}}=(\frac{\hat{\pi}_{k}}{w_{k}}\tilde{p}^{(k)}_{j})_{k\neq i,y\in\mathcal{Y}}=\Phi(Y,\mathbf{{\widetilde{p}}}_{-i})$, with $\Phi(Y,\cdot)$ coordinate-wise nondecreasing, and Assumption 4 (i) entail
$s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})=s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i})=s^{\min}_{i}(\Phi(Y,\Psi_{i}(p^{(Y_{n+i})}_{i},W_{i}))).$
Hence, $s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})$ can be written as a function $g(p^{(Y_{n+i})}_{i})$, with $g:u\mapsto s^{\min}_{i}(\Phi(Y,\Psi_{i}(u,W_{i})))$ nonincreasing and only depending on $Y$ and $W_{i}$. Applying Lemma D.1 for $c=w_{Y_{n+i}}/(m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}})$, we obtain
$\mathbb{E}_{P_{X\mid Y}}\left[\frac{{\mathbf{1}\{\tilde{p}^{(Y_{n+i})}_{i}\leq\alpha s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})w_{Y_{n+i}}/(m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}})\}}}{s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})}\>\middle|\>{\widetilde{W}}_{i}\right]\leq\alpha\frac{w_{Y_{n+i}}}{m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}}}.$
As a consequence, we derive
$\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}),P_{X\mid Y},Y)$ $\leq\alpha\sum_{k\in[K]}w_{k}\left(\frac{\sum_{i\in[m]}{\mathbf{1}\{Y_{n+i}=k\}}}{m}\frac{1}{\hat{\pi}^{{\tiny\mbox{cal}}}_{k}}\right),$
which proves (i). We deduce (ii) by a simple integration:
$\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}),P_{X,Y})$ $\leq\alpha\sum_{k\in[K]}w_{k}\mathbb{E}\left(\frac{\sum_{i\in[m]}{\mathbf{1}\{Y_{n+i}=k\}}}{1+\sum_{j\in[m]}{\mathbf{1}\{Y_{j}=k\}}}\frac{n+1}{m}\right)$
$=\alpha\sum_{k\in[K]}w_{k}(n+1)\mathbb{P}(Y_{1}=k)\mathbb{E}\left(\frac{1}{1+\sum_{j\in[m]}{\mathbf{1}\{Y_{j}=k\}}}\right)$
$\leq\alpha\sum_{k\in[K]}w_{k}(n+1)\mathbb{P}(Y_{1}=k)\frac{1}{(n+1)\mathbb{P}(Y_{1}=k)}=\alpha\sum_{k\in[K]}w_{k}=\alpha,$
by using Lemma D.2 for the last inequality.
C.4 Proof of Proposition 5.1
First define $R_{j}$ the rank of $S_{Y_{j}}(X_{j})$ in $\{S_{k}(X_{j}),k\in[K]\}$ (ordered in increasing order) for $j\in[n+m]$, and consider the slightly smaller conformal $p$-values
$\breve{p}^{(k)}_{i}=\frac{1}{n+1}\left(1+\sum_{j=1}^{n}{\mathbf{1}\{R_{j}>1\}}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{k}(X_{n+i})\}}\right)\leq\overline{p}^{(k)}_{i},\>\>i\in[m],k\in[K],$ (32)
which means that the calibration is only made with examples having a label not minimizing the score function. The rationale behind using this $p$-value rather than $\overline{p}^{(k)}_{i}$ is that, due to the post-processing, the elements $X_{n+i}$ of the test sample cannot produce an error provided that $R_{n+i}=1$ so that we can restrict the test sample to those with $R_{n+i}>1$ when computing the FCR.
Assume $K=2$. We first prove that $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})=\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}})$. It is enough to prove that the adjusted $p$-values $q_{i}$ obtained from $\mathbf{\overline{p}}$ and $\mathbf{\breve{p}}$ are the same for this non-trivial selection (since for the selected $i$, the procedure always chooses $\arg\min_{k\in[K]}S_{k}(X_{n+i})$ due to the post-processing). Letting $S_{\min}(x)=\min_{k\in[K]}S_{k}(x)$, $S_{\max}(x)=\max_{k\in[K]}S_{k}(x)$, this comes from
$\min_{k\in[K]}\{\breve{p}^{(k)}_{i}\}$ $=\frac{1}{n+1}\left(1+\sum_{j=1}^{n}{\mathbf{1}\{R_{j}>1\}}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{\max}(X_{n+i})\}}\right)$
$=\frac{1}{n+1}\left(1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{\max}(X_{n+i})\}}\right)=\min_{k\in[K]}\{\overline{p}^{(k)}_{i}\},$
where the second equality holds because $S_{Y_{j}}(X_{j})\geq S_{\max}(X_{n+i})$ is impossible for $R_{j}=1$, that is, when $S_{Y_{j}}(X_{j})$ is a minimum ($S_{\min}(X_{j})<1/2<S_{\max}(X_{j})$ almost surely by the assumptions on the score function).
Now prove Proposition 5.1, by showing the equality for $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}})$ and by carefully modifying the proof of Theorem A.1 (§ A.2) in order to get only equalities, by using that we consider the case of the non-trivial selection, that is, $\mathcal{S}=\mathrm{BH}(\mathbf{q})$, with $q_{i}=\min_{k\in[K]}\breve{p}^{(k)}_{i}$. By definition of the FCP (2), we have (remember also that the prediction set always includes $\arg\min_{k\in[K]}S_{k}(X_{n+i})$ due to the post-processing)
$\mathrm{FCP}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}}),Y)$ $=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}_{n+i}^{\alpha|\mathcal{S}|/m}(\mathbf{\breve{p}})\}}}{1\vee|\mathcal{S}|}$
$=\sum_{i\in[m]}{\mathbf{1}\{R_{n+i}>1\}}\frac{{\mathbf{1}\{i\in\mathcal{S},\breve{p}^{(Y_{n+i})}_{i}\leq\alpha|\mathcal{S}|/m\}}}{1\vee|\mathcal{S}|},$
because no error can occur when $R_{n+i}=1$. Now, since $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ and $q_{i}\leq\breve{p}^{(Y_{n+i})}_{i}$, ${\mathbf{1}\{i\in\mathcal{S},\breve{p}^{(Y_{n+i})}_{i}\leq\alpha|\mathcal{S}|/m\}}={\mathbf{1}\{\breve{p}^{(Y_{n+i})}_{i}\leq\alpha|\mathcal{S}|/m\}}$. This entails
$\mathrm{FCP}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}}),Y)=\sum_{i\in[m]}{\mathbf{1}\{R_{n+i}>1\}}\frac{{\mathbf{1}\{\breve{p}^{(Y_{n+i})}_{i}\leq\alpha|\mathcal{S}|/m\}}}{1\vee|\mathcal{S}|}.$
Let $\xi=({\mathbf{1}\{R_{j}>1\}})_{j\in[n+m]}$ and now prove
$\mathbb{E}[\mathrm{FCP}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}}),Y)\>|\>\xi]=\sum_{i\in[m]}{\mathbf{1}\{R_{n+i}>1\}}\mathbb{E}\Big{[}\frac{\lfloor(n+1)\alpha K_{i}/m\rfloor/(n^{\prime}+1)}{K_{i}}\>\Big{|}\>\xi\Big{]},$ (33)
with $n^{\prime}=\sum_{j\in[n]}{\mathbf{1}\{R_{j}>1\}}$ for some random variables $K_{i}\geq 1$, $i\in[m]$. This implies the result because the last display is at most
$\alpha\frac{(n+1)\sum_{i\in[m]}{\mathbf{1}\{R_{n+i}>1\}}}{m(\sum_{j\in[n]}{\mathbf{1}\{R_{j}>1\}}+1)}$
(with equality if $(n+1)\alpha/m$ is an integer). By Lemma D.2, the expectation of the latter is equal to $\alpha(1-(1-\mathbb{P}(R_{1}>1))^{n+1})$.
Let us now prove (33). For this, fix $i\in[m]$ with $R_{n+i}>1$ and note that for all $j\in[m]$, $j\neq i$,
$q_{j}$ $=\frac{1}{n+1}\left(1+\sum_{k=1}^{n}{\mathbf{1}\{R_{k}>1\}}{\mathbf{1}\{S_{Y_{k}}(X_{k})\geq S_{\max}(X_{n+j})\}}\right)$
$=\frac{1}{n+1}\Big{(}{\mathbf{1}\{S_{Y_{n+i}}(X_{n+i})<S_{\max}(X_{n+j})\}}+\sum_{s\in A_{i}}{\mathbf{1}\{s\geq S_{\max}(X_{n+j})\}}\Big{)}$
$=\frac{1}{n+1}\Big{(}{\mathbf{1}\{S_{\max}(X_{n+i})<S_{\max}(X_{n+j})\}}+\sum_{s\in A_{i}}{\mathbf{1}\{s\geq S_{\max}(X_{n+j})\}}\Big{)}$
$\geq\frac{1}{n+1}\sum_{s\in A_{i}}{\mathbf{1}\{s\geq S_{\max}(X_{n+j})\}}=:q^{\prime}_{j},$
by letting $A_{i}=\{S_{Y_{k}}(X_{j}),j\in[n]:R_{k}>1\}\cup\{S_{Y_{n+i}}(X_{n+i})\}$ (all distinct almost surely). The third equality above is true because $K=2$ and $R_{n+i}>1$. We apply now Lemma D.3 because $\mathbf{q}=(q_{j},1\leq j\leq m)$ and $\mathbf{q}^{\prime}=(q^{\prime}_{j},1\leq j\leq m)$ (defined as above and with $q_{i}^{\prime}=1/(n+1)$) satisfy (36). Indeed, if $q_{j}>q_{i}$ for $j\neq i,$ then $S_{\max}(X_{n+j})\leq S_{\max}(X_{n+i})$ and $q_{j}=q^{\prime}_{j}$. Hence, we obtain, by letting $K_{i}=|\mathrm{BH}(\mathbf{q^{\prime}})|$ (note that $\mathbf{q^{\prime}}$ depends on $i$)
$\{q_{i}\leq\alpha|\mathrm{BH}(\mathbf{q})|/m\}=\{q_{i}\leq\alpha K_{i}/m\}\subset\{|\mathrm{BH}(\mathbf{q})|=K_{i}\}.$
In addition, $\mathbf{q}^{\prime}=(q^{\prime}_{j},1\leq j\leq m)$ is a vector measurable w.r.t. the variable $W_{i}=(A_{i},(S_{\max}(X_{n+j}))_{j\in[m]\backslash\{i\}}).$ Now, we use that by exchangeability of the elements of $A_{i}$ conditionally on $\xi$, $W_{i}$, and $R_{n+i}>1)$ (with no ties),
$\mathbb{P}(\breve{p}^{(Y_{n+i})}_{i}\leq t\>|\>W_{i},\xi,R_{n+i}>1)=\frac{\lfloor(n+1)t\rfloor}{n^{\prime}+1}.$
Applying this with $t=\alpha K_{i}/m$ entails (33).
Appendix D Auxiliary results
Lemma D.1 (Lemma 3.2 in Blanchard and Roquain, (2008)).
Let $g:[0,1]\rightarrow(0,\infty)$ be a nonincreasing function and $U$ be a random variable which is super-uniform, that is, $\forall u\in[0,1],$ $\mathbb{P}(U\leq u)\leq u$. Then, for any $c>0$, we have
$\mathbb{E}\left[\frac{{\mathbf{1}\{U\leq cg(U)\}}}{g(U)}\right]\leq c.$
Lemma D.2 (Lemma 1 of Benjamini et al., (2006)).
If $T$ is a Binomial variable with parameter $N-1\geq 0$ and $t\in(0,1]$, we have
$\mathbb{E}[1/(T+1)]=(1-(1-t)^{N})/(Nt)\leq 1/(Nt).$
Lemma D.3 (Lemma D.6 of Marandon et al., (2024)).
Write ^=^$\widehat{\ell}=\widehat{\ell}(\mathbf{q})=|\mathrm{BH}(\mathbf{q})|$ for the number of rejections of $\mathrm{BH}(\mathbf{q})$ (23). Fix any $i\in\{1,\dots,m\}$ and consider two collections $\mathbf{q}=(q_{j},1\leq j\leq m)$ and $\mathbf{q}^{\prime}=(q^{\prime}_{j},1\leq j\leq m)$ which satisfy almost surely that
$\forall j\in\{1,\dots,m\},\left\{\begin{array}[]{cc}q^{\prime}_{j}\leq q_{j}&\mbox{ if }q_{j}\leq q_{i};\\ q^{\prime}_{j}=q_{j}&\mbox{ if }q_{j}>q_{i}.\end{array}\right.$ (36)
Then $\{q_{i}\leq\alpha\widehat{\ell}(\mathbf{q})/m\}=\{q_{i}\leq\alpha\widehat{\ell}(\mathbf{q}^{\prime})/m\}\subset\{\widehat{\ell}(\mathbf{q})=\widehat{\ell}(\mathbf{q}^{\prime})\}.$
Appendix E Additional illustrations in the regression case
E.1 Excluding a null value $y_{0}\in\mathbb{R}$
We consider here =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>y_{0}\notin I\}$, which is useful in situation where the value $y_{0}$ corresponds to some “normal” value and the user wants to report only prediction intervals for “abnormal” individuals, that is, when the outcome value deviates from this reference value. The score considered here is $S_{y}(x)=|y-\mu(x)|/\sigma(x)$, where $\mu$, $\sigma$ are predictors of the conditional mean and variance, respectively.
An illustration is provided in Figure 10 in the non-parametric Gaussian model for different situations: Rows 1,2 correspond to an homoscedastic model with perfect prediction of the variance and a mean predictor with various accuracy (less accurate in the middle for row 1, more accuracy in the middle for row 2). Rows 3,4 correspond to an heteroscedastic model with perfect prediction of the mean and a the variance predictor under-estimating the variance in the middle for row 3, and over-estimating the variance in the middle for row 4). The marginal prediction interval (no selection) is displayed in light blue while the prediction interval after selection is displayed in dark red. The quantities reported at the top of each panel are the FCR and adjusted power averaged over $1000$ repetitions (while each panel only displays the last experiment as a typical realisation of the sample).
Classical conformal InfoSP InfoSCOP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Informative selection by excluding the null value $y_{0}$ in the (homoscedastic or heteroscedastic) regression case, see § E.1.
First, as expected, the naive classical conformal selection does not provide a correct FCR control in all cases, while InfoSP and InfoSCOP do control the FCR in any case. Second, InfoSCOP always improves InfoSP in terms of power, and the range of improvement depends on where non-covering errors are likely to happen: when errors are more likely to arise on the selection (rows 1,3), the behavior of InfoSP and InfoSCOP are similar; when errors are less likely to arise on the selection (rows 2,4), InfoSCOP improves over InfoSP in a striking manner (and is even better than classical conformal). This is both because of the reduction of the selection effect and because the pre-processed $p$-values are calibrated with much lower residuals and thus are much more efficient than the original $p$-values.
E.2 Prediction intervals for Jin and Candes, (2023) selection
We focus here on the case where the user wants to build prediction sets only for outcomes $Y_{n+i}>y_{0}$, which corresponds to excluding the set $\mathcal{Y}_{0}=(-\infty,y_{0}]$ and is related to the selection proposed in Jin and Candes, (2023).
We consider here the aim of finding one-sided prediction intervals on the selected. Let us assume that the score function is monotone in the following sense (Jin and Candes,, 2023): for all $x\in\mathbb{R}^{d}$, for $y\leq y^{\prime}$, $S_{y}(x)\geq S_{y^{\prime}}(x)$. Examples of monotonic scores are given by $S_{y}(x)=(\mu(x)-y)/\sigma(x)$, $S_{y}(x)=(\mu(x)-My)/\sigma(x)$ with $M=2\sup_{x}\mu(x)$. The following result summarizes our finding in this case.
Corollary E.1.
Consider the iid model in the regression case and assume that the score function is monotone (see above) and such that Assumption 2 (ii) (iii) and Assumption 1 hold. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap(-\infty,y_{0}]=\emptyset\}$ and $p$-value collection $\mathbf{\overline{p}}$ (7):
  • (i)
    InfoSP selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ the rejected set of BH procedure at level $\alpha$ applied with the $p$-values
    $q_{i}=\overline{p}^{(y_{0})}_{i}=\frac{1}{n+1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{y_{0}}(X_{n+i})\}}\Big{)}.$
  • (ii)
    The selection $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ of InfoSP controls the FDR at level $\alpha$ in the following sense:
    $\sup_{P_{XY}}\mathbb{E}_{(X,Y)\sim P_{XY}}\left[\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\leq y_{0}\}}}{1\vee|\mathcal{S}|}\right]\leq\alpha.$
  • (iii)
    The prediction intervals are of the form $\mathcal{C}_{n+i}=\{y>y_{0}\>:\>S_{y}(X_{n+i})\leq S_{(n_{\alpha})}\}$, $i\in\mathcal{S},$ where $S_{(1)}\leq\dots\leq S_{(n)}$ are the ordered calibration scores $S_{Y_{j}}(X_{j})$, $1\leq j\leq n$ (with $S_{(n+1)}=+\infty$), and $n_{\alpha}=\lceil(1-\alpha|\mathcal{S}(\mathbf{p})|/m)(n+1)\rceil$.
  • (iv)
    These prediction intervals control the FCR at level $\alpha$ in the sense of (17).
Importantly, (iii) ensures that the produced prediction intervals are always “doing what they should do”, by always excluding values smaller than $y_{0}$.
Proof.
The proof is direct with Example 2.1 and monotonicity (for (i)), Lemma 2.1 (for (ii)), Remark 2.1 (for (iii)), Theorem 3.1 (for (iv)). ∎
Obviously, a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3. Corollary E.1 is illustrated on Figure 11 when $\mathcal{S}^{(0)}$ is taken here has $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ with the score $S_{y}(x)=(\mu(x)-y)/\sigma(x)$. The comments are qualitatively similar to those of Figure 10: when covering errors are less likely on the selection, the improvement of InfoSCOP over InfoSP is substantial.
Classical conformal InfoSP InfoSCOP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Same as Figure 10 with Jin and Candes, (2023) type selection and one-sided prediction intervals, see § E.2.
E.3 Comparison with existing selective prediction sets
Figure 12: Comparison to previous conformal selective inferences in the regression case: new informative prediction intervals (InfoSCOP) versus non-informative prediction intervals (SCOPab) (left: homoscedastic Gaussian regression with perfect variance prediction and errors in the mean prediction; right: heteroscedastic Gaussian regression with perfect mean prediction and errors in the variance prediction). Informative means here prediction intervals that does not contain $y_{0}$ (dashed line), see § E.3.
Figure 12 is useful to visualize the difference between the selection proposed by Bao et al., (2024) and our informative selection. We display selective prediction intervals for two FCR controlling selections proposed in Bao et al., (2024): SCOPa uses a thresholding rule $\mathcal{S}=\{i\in[m]\>:\>\mu(X_{n+i})\geq y_{0}\}$, while SCOPb selects the largest $\mu(X_{n+i})$’s. Each time, a suitable conditional conformal prediction set is built in Bao et al., (2024) and reported in Figure 12 in green and purple. As one can see, either the selection is too conservative, or the prediction interval could include the nominal $b_{0}$. This not the case of InfoSCOP which always exclude $y_{0}$ by essence.
Appendix F Additional illustrations in the classification case
F.1 Three classes, each a mixture of bivariate normals
For the iid model, as in § 5.3, we consider the balanced and unbalanced cases of 3 classes. However, here we consider that each class comes from a mixture of two bivariate normals, where one component of the mixture is identical in all three classes, and given the class there is probability half of coming from that common mixture component. Thus the overlap between the classes is much greater than in the settings considered in § 5.3.
In Figure 13 we show that the qualitative conclusions with regard to the respective procedures are as in § 5.3. Interestingly, in the hardest setting at the bottom row we see that the power of InfoSCOP is even larger than classic conformal. This is because the initial selection step not only prevents paying too much for selection, but it can also improve the calibration set for computing the conformal $p$-values after selection. In this hard setting, there is a large improvement of the tail probability that matters for inference. We demonstrate this in one realization of the data generation in Figure 14, where we can see that the 0.95 quantile of the original calibration set is much larger than the 0.95 quantile of examples that remain in the calibration set after the initial selection step. This is because the data generation is such that examples in the central cloud tend to have larger nonconformity scores $S_{Y_{i}}(X_{i})$ than the examples that are not in the central cloud, and most of the examples from the central cloud were eliminated from the calibration sample with the initial selection step. So the remaining nonconformity scores after the initial selection step tend to be smaller. Since the nonconformity scores of the examples from the test sample are unchanged, the improved calibration set after the initial selection step results in $p$-values that tend to be smaller when testing the null group, and in particular there are many more $p$-value that are below the 0.05 level.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Selecting informative prediction sets in the iid setting with a large overlap between three classes. FCR (left column), resolution-adjusted power (middle column) and the expected fraction of covering prediction sets (right column) versus SNR for: balanced classes for minimally informative prediction sets (first row) and for prediction sets excluding a null class (second row); unbalanced classes for minimally informative prediction sets (third row) and for prediction sets excluding the largest class (fourth row). Each class is a mixture of a common component and a unique component, see § F.1 for details. Based on 2000 data generations, 1000 data points were used for training, and $n=m=500$.
Figure 14: The setting is that of unbalanced classes, with each class having probability half of being in the same mixture component, and the SNR is 8. In the first row, the data points from each of the three classes, where the null group is in blue: top left panel for the entire test sample, top right panel the remaining examples after the initial selection step. In the second row, left panel, their $p$-values using the entire calibration set (black circles) and using the examples from the calibration set remaining after the initial selection step (blue triangles). In the second row, right panel, the empirical CDF of the nonconformity scores for true classes in the entire calibration set (gray) and using the examples from the calibration set remaining after the initial selection step for the of the (blue).